A Complete Bibliography of the LMS Journal of Computation and Mathematics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 January 2018 Version 2.15

Title word cross-reference

#P [20]. (2 × 2 · G): 2 [160]. 1 [174]. 12 [190]. 13A [321]. 2 [154, 269, 216, 117, 317, 146, 172, 6, 51]. 3 [216, 166]. 30 [11]. 4 [124]. 47: 23 [341]. 5 [332]. 6 [276, 142]. 6560 [54]. 7 [354]. 8 [267]. 2F_{4}(q^{2}) [143, 161]. a^{2} − 2 = c^{p} [211]. a^{2} − 2b^{6} = c^{p} [211]. A_{n} [85]. B [73]. C^{3} [264]. C_{3,4} [122]. d [198]. E_{6} [123]. E_{7}(2) [306]. E_{8} [123, 147]. ℓ [295]. F_{16} [126]. Fi′_{24} [132]. GF(2) [104]. GF(p^{6}) [369]. GL_{2}(Q_{p}) [249]. j [152]. J_{4} [196, 59]. K [148, 230]. K3 [271]. K^{2} = 3 [332]. L [263, 362, 189, 290]. λ [323]. M [173]. M_{22} [129]. M_{24} [146]. F_{2} [347]. F_{p^{n}} [281]. Q [289]. Z_{p}[[u]] [268]. PSL(2, p) [248]. PSL_{2}(27) [251]. p [73, 173, 145, 186, 286, 340, 249, 255, 130, 346, 194]. p_{g} = 0 [332]. ±2^{a}3^{b} [85]. (mod p) [239]. PSL(2, R) [151]. PSL_{2}(13) [321]. Q [355]. S_{n} [85]. SL_{2} [126]. SL_{2}(R) [269]. θ [206, 243]. U_{7} [209]. W [174]. X_{0}(125) [68]. X_{0}(5) [87]. X_{0}(N) [164]. Y_{n}(p^{a}) [346]. Z^{2} [96].

-Adic [130, 186, 286, 340, 249, 194]. -algebras [174]. -Bessel [230]. -complete [20]. -coverings [276, 51]. -curve [355]. -Cycles [142]. -dimensional [174]. -elements [321]. -Engel [124]. -functions [263, 189, 290]. -Groups [148]. -Harish-Chandra [198]. -homogeneous [6]. -Invariant [152]. -invariants [323]. -isogeny [295]. -Local [146]. -Maruyama [243, 206]. -Regular [145]. -series [362]. -subgroups [255, 346, 172]. -torsion [332].

132 [56].

2 [34]. 2000f [34]. 2B [128].

6 [50].

A. [191]. abelian [218, 309, 279, 272, 199]. Abelianization [242]. above [358]. absolute [258]. Accelerating [212]. acceptance [18]. Account [111]. Accurately [105]. acting [208, 209]. Action [128, 147, 301]. actions [6, 8]. addition [218]. Adic [130, 186, 286, 340, 249, 194]. adjoint [159, 4]. Adomian [373]. after [21]. against [350]. Airy [260]. alcove [63]. algebra [268, 304, 58]. Algebraic [148, 141, 120, 99, 199]. Algebras [120, 135, 104, 138, 359, 308, 215, 174, 16, 178, 63, 223, 292, 342]. Algorithm [98, 153, 103, 93, 273, 282, 364, 19, 330, 317, 325, 167, 292, 281, 58]. Algorithmic [79, 224, 6, 8]. Algorithms [351, 371, 230, 7]. Almost [206, 243]. Alperin [59, 73, 132, 173]. AMD [7]. AMD-K7^{T M} [7]. ameliorated [345]. among [9]. Analysis [79, 47, 345, 28, 206, 243, 23, 187, 22]. analytic [340, 200]. Anti [96]. Anti-Ferromagnetic [96]. any [365]. aperiodic [297]. Application [167, 348, 340, 368, 238, 166]. applications [371, 277, 70]. applied [206, 243]. approach [318, 328, 25, 35, 367]. Approximate [127, 351]. Approximating [170, 275]. Approximation [245, 118, 262, 167, 241]. Approximations [260]. arbitrary [158]. arguments [335, 320]. arising [16, 112]. Arithmetic [147, 253]. array [339]. article [243]. Artin [43]. aspects [24]. Associated [87, 199, 21, 174]. association [256]. asymptotic [206, 243]. Atkin [305]. attached [323, 290]. Attack [74, 47, 350]. Automating [72]. Automorphic [140, 249, 193]. automorphism [216]. Automorphisms [164]. auxiliary [328]. average [285, 305]. axiom [62].

B [229]. B-splines [229]. Baby [73, 341, 128, 48]. Barker [250]. barycentres [75]. Base [103]. Based [99, 273, 356, 44, 373, 167, 374]. bases [224, 270, 35]. basis [296, 162, 58]. Baumslag [182]. BDD [44]. BDD-based [44]. Bernoulli [163]. Bespoke [307]. Bessel [230, 260, 265]. beta [267]. between [12, 70]. BGJT [282]. BGJT-algorithm [282]. Bianchi [242, 236]. binary [10, 34]. binomials [69]. bipartite [183]. Birch [200]. Black [92, 61, 172]. Black-Box [92, 61, 172]. blob [63]. blocks [342]. Borcherds [237]. Borel [143]. Bound [97, 163]. Boundary [108]. bounded [278, 241]. Bounds [230, 113, 83, 96, 252, 261, 69, 27]. Box [92, 225, 61, 172]. branched [57]. branching [167]. Brauer [125, 216, 42]. breaking [72]. Brownian [344]. Bruhat [249]. Buchberger [162]. buildings [208].

Calculating [255]. calculation [287, 44]. Calculations [140, 72]. Calculus [150, 14, 314]. canonical [249, 372, 254, 58]. capture [55]. cardinalities [355]. case [66]. cases [198]. Cassels [179]. Categorical [116]. census [347, 289]. central [246]. certain [343, 3, 163, 170, 16, 55]. Chamber [146]. chambers [214]. Champ [304]. champion [231]. Chandra [198]. Character [115, 143, 161, 367, 160, 191]. Characteristic [99, 117, 283, 365, 33, 49, 164, 36, 47, 270, 158]. characteristics [216]. characters [246, 311, 346, 216]. checked [7]. Cherednik [304]. Chevalley [255, 346, 36, 224]. Chief [136]. Chlodowsky [245]. choice [62]. chromatic [217]. ciphers [220]. Class [91, 287, 294, 128, 293, 371, 186, 340, 368, 235, 48]. Classes [151, 255, 55]. Classical [84, 110, 318, 194]. Classification [251, 157]. classifying [24]. Closures [114]. CM [366, 313, 212]. codes [231]. coefficients [261, 343, 204, 263, 65]. Cohomology [119, 153, 312, 334, 181]. Coleman [303]. Collecting [369]. collection [2]. colour [177]. Colours [149]. Com [330]. Com-Poisson [330]. combinations [44]. combinatorial [157, 71]. committed [53]. common [351]. commutativity [203]. Compact [153, 322]. Complete [218, 20, 155]. completely [375]. completeness [188]. Completing [42]. Complex [229, 4, 348, 178, 188]. Complexity [234, 177]. compliance [7]. composite [47, 294, 222]. composition [245]. comprehensive [226]. Compression [127]. Comput [34]. Computation [257, 114, 94, 130, 270, 293, 156, 13, 3, 207, 284, 345]. Computational [24, 326]. Computations [148, 194, 368, 166]. compute [58]. Computing [191, 90, 239, 141, 315, 246, 67, 101, 213, 291, 249, 237, 327, 285, 362, 136, 334, 358, 43, 106, 300, 355, 372, 137, 361, 41, 227, 302, 272, 138, 371, 215, 317]. concentration [17]. concerning [348]. Condensation [210, 223]. Condensed [135]. Condition [108, 185]. conductor [294]. Congruence [151, 259, 168]. Conjectural [56]. Conjecture [132, 59, 73, 173, 253, 200, 132]. Conjugacy [91, 151, 128, 48, 255]. conservation [307]. consistency [62]. Constructing [366, 274, 279, 12, 84, 165, 178]. Construction [112, 301, 123, 224]. constructive [23, 61, 187, 269]. contains [321]. continuity [66, 228, 187]. contribution [40]. Control [109, 116]. Convection [121]. Convection-Diffusion [121]. Convergence [131, 118, 345]. convex [23]. convolutions [163]. Conway [81]. coprime [205]. correlated [330]. correspondences [349]. Corrigenda [46]. Corrigendum [243, 34]. Coset [35]. Counting [153, 20, 117, 348]. coupled [374]. Coupling [109, 45]. Covariance [112]. Covering [88]. coverings [276, 57, 51]. covers [78]. Coxeter [215, 327]. CoxIter [327]. criteria [162]. criterion [343]. critically [289]. Cross [77]. crossed [9]. cryptanalysis [364]. Cryptography [99, 280, 279]. cubature [337]. cubic [10, 34, 323, 144]. cup [334]. Curve [74, 280, 354, 355]. Curves [94, 86, 117, 134, 139, 93, 122, 303, 370, 324, 363, 313, 299, 171, 316, 274, 266, 199, 276, 249, 12, 285, 362, 353, 336, 164, 43, 277, 200, 270, 372, 70, 301, 166, 305, 41, 192, 51, 219]. Cuspidal [332, 37, 5, 236]. Cut [97]. Cycles [142]. Cyclic [114, 323, 190]. cyclic- [190]. cyclotomic [204, 294].

Dade [59, 132]. Darboux [39]. data [330]. database [363, 38]. Databases [370]. decay [67]. deciding [33]. decision [76, 162]. Decomposing [133]. Decomposition [198, 373, 190, 254]. deduction [44]. Deep [2]. defined [313, 289]. defining [339, 368, 36]. Definite [140]. degenerate [201]. Degenerations [181]. degree [303, 293, 54, 267, 36, 47]. Degrees [88]. del [157]. Delay [131, 28, 241]. Dem’Janenko [94]. d’enfants [240]. densest [275]. density [176]. Dependencies [120]. derivatives [167]. descent [47]. Deslauriers [261]. dessins [240]. destabilising [206, 243]. determinant [284]. Deterministic [256]. Deutsch [98]. DHP [70]. Diagrams [89]. diamonds [31]. difference [307]. Differential [131, 118, 105, 28, 335, 64, 16, 373, 60, 320, 374, 326, 241]. Diffie [76]. Diffusion [121, 109]. Dimension [201, 104, 16, 317]. dimensional [11, 174, 32, 46]. Diophantine [87, 266].

Dirac [156]. directed [31, 183]. Dirichlet [343, 52, 189]. Discovering [183]. Discrete [82, 74, 150, 243, 278, 282, 328, 269]. Discriminant [85]. discriminants [298, 40]. disk [254]. dispersion [330]. distribution [305]. distributions [370, 356]. divisibility [30]. Division [359, 7]. divisor [351]. Divisors [144]. DLP [70]. Domain [111, 25]. Domain-Theoretic [111, 25]. domains [168, 249]. double [344, 262]. dream [319]. drift [344]. Drinfeld [324]. Duality [150]. Dubuc [261]. Durrmeyer [245]. Dyer [200]. dynamics [15].

Easy [76]. ECDLP [47]. ECM [353]. Effective [352]. Efficient [130, 221, 197, 78]. eigenstates [156]. eigensystems [239]. Eigenvalue [159]. eigenvalues [52]. Elements [145, 354, 321]. Elkies [305]. Elliptic [74, 108, 134, 370, 280, 316, 274, 266, 30, 276, 354, 12, 240, 175, 336, 200, 270, 372, 70, 305, 41, 51, 219]. embedding [249]. enclosures [159, 4]. encoding [364]. endomorphism [274]. Endpoints [97]. Engel [124]. engine [50]. entropy [163]. Enumerating [57]. enumeration [35]. enveloping [58]. Equation [87, 375, 318]. Equations [131, 118, 105, 28, 17, 335, 211, 64, 16, 373, 60, 320, 374, 326, 241]. equivalence [18, 70]. Ergo [50]. error [356, 260, 53]. estimates [53]. estimating [330, 226]. Estimation [75, 308, 180]. eta [247]. eta-quotients [247]. Euclidean [29]. Euler [131]. Evaluating [263]. evaluation [371]. even [303]. even-degree [303]. Every [321]. Evidence [56]. Exact [127, 190]. Examples [225, 313, 271]. Exceptional [138, 198, 174, 342]. exclosures [159]. existence [171, 187]. expansion [220, 267]. expansions [155]. experimental [35]. Explicit [365, 27, 140, 158, 238, 166, 28, 260, 349, 40]. explicitly [192]. Exponential [64]. exponentials [188]. exponents [278].

Extending [98]. extension [162].

Extensions [114, 38]. exterior [19]. Extra [86].

factoring [256]. factorisation [71].

factorization [232]. factorizations [234]. factors [65]. fairness [185]. Faithful [88]. Families [117, 192, 324]. Fast [371, 61, 122]. fat [333]. Fermat [144]. Ferromagnetic [96]. few [263, 227]. fibering [254]. Field [136, 122, 283, 313, 369, 368, 336, 360, 38]. Fields [114, 134, 85, 283, 234, 293, 45, 316, 323, 33, 12, 49, 233, 43, 277, 47, 294, 355, 41]. filtering [197]. Finding [341, 120, 281, 325]. Finite [145, 98, 136, 296, 122, 283, 5, 213, 12, 215, 174, 255, 307, 49, 360, 43, 277, 36, 289, 47, 355, 195, 184]. finiteness [33]. Finsler [202]. first [52]. Fischer [48]. Five [149]. flag [312]. floating [7]. form [243]. formalisation [14]. Forms [107, 140, 297, 10, 34, 249, 209, 288, 194, 232, 236, 290, 361, 284, 302]. formula [243, 340, 221, 40]. formulas [158]. four [177, 248]. fourth [337]. fourth-order [337]. fractional [374]. fractional-order [374]. Free [9, 297, 343, 329, 269]. friendly [309]. Frobenius [338, 181]. full [242]. Function [114, 134, 230, 265, 331, 241]. functional [60]. Functions [107, 106, 133, 375, 315, 187, 199, 263, 229, 287, 347, 260, 358, 43, 258, 289, 262, 189, 290, 227]. functors [3]. Fundamental [168, 249].

G [191]. Galois [126, 257, 85, 217, 147, 272]. Garsia [163]. genera [235]. general [4, 5]. Generalised [92]. Generalising [74]. Generalization [56]. Generalized [63, 374]. Generate [93]. generated [288]. Generating [119, 248]. generic [50]. Genus [154, 299, 94, 366, 363, 313, 168, 199, 362, 358, 166, 254]. genus- [166]. Genus-2 [299, 363]. Genus-3 [94, 366]. genus-one [168]. genus-zero [168]. Geometric [93, 16, 314]. geometrically [362]. geometricity [310]. Geometry [146, 202, 63]. GGH [364]. GHS [74, 47]. given [359, 299, 274, 338]. Global [150, 53]. Good [324]. graded [16]. Graph [127, 146]. graphs [176]. Gröbner [162, 35]. Group [132, 81, 129, 136, 85, 122, 59, 339, 160, 293, 371, 26, 207, 354, 368, 259, 216, 42, 203, 196, 8, 91, 128]. Groups [92, 145, 98, 82, 148, 126, 113, 208, 141, 101, 143, 84, 88, 103, 140, 124, 110, 80, 370, 26, 191, 61, 9, 78, 1, 246, 5, 33, 54, 213, 11, 257, 182, 76, 215, 255, 327, 170, 161, 311, 346, 32, 46, 165, 36, 178, 217, 195, 301, 242, 184, 172, 6, 367]. Growth [182].

Hall [213]. Halley [345]. Harada [216]. Hardy [221]. Harish [198]. harmonic [17]. Hasse [285, 352]. Hausdorff [25]. having [262]. Hecke [239, 215, 178]. height [370, 278, 175]. Heights [130, 372]. Hellman [76]. Hermite [27]. Hessian [353]. heuristic [371]. Heuristics [309]. hexagons [337]. hidden [169]. High [266, 283, 214]. high-characteristic [283]. High-rank [266]. Higher [83, 242]. Higher-Order [83]. Hilbert [56, 119, 334, 139, 290]. HJM [167]. Homaloidal [333]. Homogeneous [95, 6]. homogenization [201]. homomorphism [210]. Homomorphisms [127, 102]. homotopical [3]. homotopical-functors [3]. Høyer [98]. Hurwitz [229, 110, 80]. hybrid [197]. hydrodynamics [159]. Hyper [280]. Hyper-and-elliptic-curve [280]. hyperbolic [208, 327]. Hyperelliptic [316, 86, 117, 139, 303, 366, 285, 362, 164]. hypersurfaces [227].

ideal [292]. ideals [244, 333]. Identifying [219]. identities [9]. IEEE [7]. II [250]. III [125]. imaginary [230]. implementation [49, 353, 221]. implicit [187]. Improved [113, 96]. incremental [322]. Indefinite [121, 292]. independent [308]. induced [266]. inequality [275]. infinitely [170]. inputs [328]. integer [339]. integers [291, 205]. Integral [129, 120, 114, 26, 334, 235]. integration [303, 25, 16]. Interactions [105]. interest [167]. interest-rate [167]. Interference [116]. Intermediate [80]. Interpolating [93, 155]. interpolation [228]. Interpolations [102]. interpolatory [261]. Intersection [79, 352]. intervals [205]. Invariant [152]. invariants [359, 186, 279, 323, 327]. inverses [204]. Inversion [112, 345]. Involutions [81, 86, 196]. Involutive [95]. Irreducible [123, 190, 58]. Isabelle [62]. Isabelle/ZF [62]. isogenies [316, 365, 12, 300]. isogenous [192]. isogeny [371, 295]. isomorphic [251]. Isospectral [107]. Itô [243]. Iteration [142]. iterative [326]. Iwahori [215].

J [34]. Jacobi [361, 374]. Jacobian [122]. Jacobians [366, 299, 315, 192]. Janko [196]. JKL [353]. JKL-ECM [353]. Jordan [342]. Jozsa [98].

K3 [348, 240, 334, 347]. K7^{T M} [7]. Kagome [149]. key [350]. known [263]. Kolmogorov [109]. Kravchuk [252]. Kummer [207, 158]. Kusuoka [167]. Kutta [53, 314].

Lagrange [228]. Laplace [344]. Laplacian [52]. Large [103, 122, 293]. Large-Base [103]. Largest [81]. Lattice [149, 357]. Lattices [147, 228, 235]. laws [218, 307]. Lehmer [21]. Lens [107]. less [345]. level [364, 239, 236]. Li [343]. Li-type [343]. Lie [120, 174, 16, 301, 342, 104, 138]. limit [181]. Linear [268, 120, 339, 225, 231, 5, 358, 317, 372, 184, 222]. LMS [34]. Local [146, 234]. Locus [97]. log [157]. Logarithm [74, 150, 278, 328]. Logarithms [82, 282]. logically [55]. longitudinally [330]. loops [193]. Lopatinskij [108]. Low [32, 46, 104, 364, 175, 40]. Low-dimensional [32, 46]. low-level [364]. lower [163]. LWE [356]. Lyons [42].

M. [191]. Magma [304]. magnitude [261, 204]. magnitudes [226]. Manifolds [93, 312, 301]. Manin [125, 94]. map [364]. Markov [169, 67]. Maruyama [243, 206]. mask [261]. Matching [135, 223]. Math [34]. Mathieu [129]. Matrices [127, 137, 198, 285]. Matrix [101, 136, 112, 308, 33]. Maximal [115, 84, 146, 147, 306, 359, 161, 165, 258]. McCammond’s [297]. mean [64]. mean-square [64]. Means [100, 202]. measurable [241]. mechanically [7]. mechanized [62]. Mechanizing [22]. median [180]. Medians [202]. medium [283]. medium- [283]. meet [350]. meet-in-the-middle [350]. Meixner [252]. membership [269]. memory [350]. Method [94, 106, 15, 206, 243, 45, 373, 314, 212, 326]. methods [28, 53, 345]. Microscale [105]. middle [350]. Minimal [95, 276, 88, 233, 137, 254, 51, 176, 248, 40, 342]. Mixing [96, 149]. mixture [169]. Model [96, 68, 167]. Models [116, 105, 303, 169, 276, 51]. Modular [154, 90, 240, 336, 324, 316, 204, 259, 209, 164, 194, 317, 236, 290, 37, 284, 272]. module [58]. Modules [135, 9, 5, 223, 342]. moduli [277]. monomials [227]. monotone [335, 320]. monotonicity [258]. Monster [73, 128, 173, 91, 115, 341, 48, 251, 321]. Mordell [318, 171, 270]. Mori [319]. most [54, 11]. motion [344]. MR [34]. Muller [185]. multi [225, 374]. multi-box [225]. multi-term [374]. multilinear [364]. multiple [283, 247, 307]. multiplication [271, 348, 349, 7]. multiplicative [314]. multiplier [170]. multiset [19]. multitransition [185]. Multitude [105]. multivariate [228]. natural [228]. nets [333]. networks [183]. Nice [78]. Nilpotent [308, 101, 138, 170, 184]. Nilpotent-independent [308]. Non [156, 110, 159, 171, 335, 330, 320, 264]. non-existence [171]. Non-Hurwitz [110]. non-monotone [335, 320]. non-self-adjoint [159]. non-stationary [330, 264]. Non-variational [156]. noncyclotomic [65]. nondegenerate [227]. nonic [233]. nonlinear [373, 53, 197, 326]. nonstandard [22]. Nonvanishing [290]. normal [297]. normalisation [14]. Normalizers [172]. norms [69]. Norton [216]. Note [40, 261, 262, 195]. NTRU [364, 350]. Number [145, 85, 97, 283, 214, 293, 299, 323, 329, 340, 287, 369, 368, 41]. Numbers [120, 363, 338, 267, 294, 222]. Numerical [28, 56, 83, 338, 60, 190, 23, 64, 374]. Numerical-symbolic [190].

Obstruction [125, 166]. odd [73, 173, 259]. OM [234]. one [28, 239, 168, 200, 236]. one-step [28]. ones [57]. Operations [122]. Operator [121, 209]. operators [156, 4, 13, 245, 201]. Optimal [109, 300, 166]. Optimally [109]. orbitals [8]. Orbits [138, 77]. Order [83, 371, 230, 11, 354, 296, 337, 373, 374]. ordered [370]. orders [359]. ordinary [270, 326]. orthogonal [339, 311, 165]. Oscillation [335, 320]. oscillator [17, 53]. oscillatory [27]. overconvergent [209, 302]. overlattices [273].

package [319, 304]. Pairing [99, 309, 179]. Pairing-Based [99]. pairing-friendly [309]. pairs [250, 6]. pairwise [205]. parabolic [161]. parallel [325]. parameters [330]. Parametrizing [277]. Part [201]. Partial [105, 331, 374]. path [295]. Pathwise [118]. Perfect [169]. Permutation [103, 339, 1, 54]. permutations [14]. perturbation [226]. perturbations [206, 243]. perturbed [17, 197]. Pezzo [157]. Phylogenetic [112]. Picard [214, 82, 111, 207]. planar [244, 52]. plane [277]. Point [153, 348, 117, 244, 337, 7]. Points [119, 94, 175, 299, 171, 334, 333]. Poisson [330]. pole [262]. Polynomial [126, 102, 256, 285, 55]. polynomial-time [55]. Polynomials [154, 90, 65, 137, 142, 234, 204, 21, 257, 27, 368, 252, 360, 317, 217, 226, 374]. positive [33, 164, 235]. post [289]. post-critically [289]. posterior [169]. potentials [156]. Potts [96]. power [203]. powers [244]. practical [70]. precision [286]. prefix [35]. prescribed [65]. presentation [312]. presentations [9, 78, 11]. presented [170]. preserve [307]. pricing [167]. prime [329, 322]. Primes [21, 30, 247, 305, 302]. Primitive [144, 103, 54]. Primitivity [184]. principal [292]. principle [352]. priori [53]. private [350]. Problem [74, 150, 112, 278, 328, 269, 351, 295, 296, 71, 292, 139]. Problems [109, 108, 100, 18, 159, 364, 71]. processor [7]. product [195]. Products [205, 69, 237, 334]. program [339, 55]. Programming [44, 339]. progressions [253]. Prolog [50]. Promise [98]. proof [23, 7, 50]. proofs [14]. Property [79, 195]. Proving [200, 171]. pseudo [338]. pseudo-Frobenius [338]. PyCox [215].

Quadratic [142, 371, 316, 365, 232, 289]. quadrics [352]. quantized [58]. quartic [10, 34, 347]. quartics [277]. quasi [32, 46, 358, 300, 317, 372]. quasi-linear [358, 317, 372]. quasi-optimal [300]. quasi-simple [32, 46]. quaternion [295, 292]. quintics [332]. quotients [315, 247, 291, 170].

Rademacher [221]. radially [156]. Radical [136]. radius [258]. Ramanujan [221]. ramified [247]. random [45, 201]. randomization [203]. Range [52, 23]. Ranges [83]. Rank [80, 198, 266, 269, 255, 200, 41]. Rankin [275]. Ranks [134, 370]. Rapid [149]. rare [310]. rate [167]. Rational [94, 258, 142, 363, 171, 257, 289]. rationality [182]. rationals [38]. Real [349, 343, 271, 348, 22, 294]. recognising [19]. recognition [61, 6, 8]. recurrence [222]. Recursive [106]. Reduced [350]. Reducing [368]. Reduction [10, 34, 103, 201]. reductions [355, 305]. Reductive [141]. Ree [143, 161]. Refined [82]. reflection [178]. reflex [313]. region [27, 265]. regions [343]. regression [330]. Regular [145, 337]. Related [109, 268, 102]. relations [371, 9, 369]. Relative [148, 62]. relaxation [339]. Replicable [133, 375]. Representation [31]. Representation-directed [31]. Representations [88, 174, 32, 46, 36, 178, 272]. Representatives [91, 48]. resolutions [9]. Resolving [105]. resonances [4]. respect [14]. restricted [65]. Restricting [311]. Restrictive [18]. result [39]. resultants [281]. retarded [335, 320]. reversible [67]. Revisited [116, 297]. Riemannian [75, 100, 93, 180]. rigid [181]. Ring [129, 356, 274, 312]. ring-based [356]. rings [26, 268, 291]. robust [330]. root [325, 7]. root-finding [325]. Roots [360, 336, 226, 281]. Rosenhain [279]. Rudin [253]. Runge [53, 314].

Salem [267]. sampling [155]. Sato [367]. Scheme [131, 261, 374, 264]. Schemes [119, 118, 256, 307, 334, 55]. Schreier [43]. Schrödinger [13]. Schubert [312]. Schur [170]. Schwarzian [375]. Searching [193]. self [159, 4]. self-adjoint [4]. Selmer [370]. semigroup [67]. semigroups [297, 338, 296]. semilocal [345]. separable [300]. separated [192]. sequence [222]. sequences [250, 21, 30, 155]. sequent [14]. Series [119, 136, 343, 198, 220, 362]. Serre [166]. set [338]. Sets [94, 208, 308, 248, 188]. Seven [231, 255, 337]. seven-point [337]. Seventeenth [139]. Shafarevich [354]. shape [160]. Sharpiro [108]. shifted [374]. Shimura [249]. short [205]. Siegel [287, 284]. sieve [283, 273, 171, 369]. sieves [322]. sieving [357]. Sifting [92]. Signature [150]. Simerka [232]. Simple [132, 145, 81, 129, 88, 128, 135, 104, 138, 59, 78, 67, 174, 32, 46, 216, 193, 42, 223, 196]. simply [208]. simulation [169, 374]. Single [235]. Single-class [235]. Singular [247]. singularities [24]. singularly [197]. six [255, 296]. size [248]. Slopes [209]. Small [99, 36, 78, 193, 302]. Smooth [93, 277, 301]. Software [49, 319]. Solitar [182]. Soluble [136]. Solutions [100, 64, 60]. solvable [54, 233]. solving [326]. Some [123]. space [66, 209, 277]. Spaces [107, 319, 25, 245, 210]. sparse [360]. Spatial [149]. special [311]. specific [200]. Spectral [17, 4, 83]. Spider [89]. spline [301]. splines [225, 229]. Sporadic [113, 88, 128, 26, 216, 42]. Spread [113]. square [329, 19, 64, 7]. square-free [329]. Squarefree [298]. squares [253]. stabilising [206, 243]. stability [206, 243, 64]. stabilizers [342]. Stable [68]. stationary [330, 264]. step [28]. Stochastic [131, 118, 105, 28, 206, 243, 64, 60, 197, 241]. stream [220]. Strong [66, 149]. structure [26]. structures [181, 55]. sub [67]. sub-Markov [67]. subdivision [261, 264]. Subexponential [293]. Subgroup [115, 143]. Subgroups [151, 84, 123, 306, 168, 246, 213, 269, 182, 255, 259, 240, 161, 346, 165, 172, 251]. sublattice [275]. subsets [195]. substructure [183]. subtrees [20]. successive [281]. suffices [18]. sum [329]. sums [331]. super [345]. super-Halley [345]. supersingular [274, 219]. Support [153]. sure [206, 243]. surface [57, 158]. Surfaces [119, 214, 279, 207, 271, 348, 240, 175, 334, 157, 347, 332, 254, 272]. Swinnerton [200]. Swinnerton-Dyer [200]. Sylow [255, 346]. Symbolic [2, 244, 44, 190]. symbols [37]. symmetric [156, 264]. symmetry [339, 72]. symmetry-breaking [72]. Syntactic [116]. system [206, 243, 197]. systems [185, 374]. Százs [245].

Table [115, 143, 160]. tables [191, 161]. Tate [179, 354, 367]. technology [50]. tensor [71]. tensor-factorisation [71]. term [260, 374]. test [206, 243]. Testing [203, 259, 184]. their [78, 315, 204, 88]. Theorem [56, 177, 244, 226, 238, 111]. Theoretic [111, 25]. theory [367]. Theta [15, 358]. third [373]. third-order [373]. Thought [2]. Three [11, 199, 49, 362, 205, 52]. Three-dimensional [11]. thresholds [13]. time [365, 285, 358, 300, 317, 372, 55]. Tits [249]. Tori [99]. toric [157]. Torsion [129, 266, 175, 242, 332]. Tracking [286]. Transfer [127]. transform [344]. transition [265]. transitive [1]. transitively [208]. transportable [37]. Traps [282]. Treatment [373]. tree [249, 20, 167]. tree-based [167]. Trees [112, 42]. triangles [176]. triangular [208]. trinomial [298]. tripartite [176]. triple [195]. triples [266]. Tropical [152]. truncated [344]. Tuple [357]. twelfth [336]. Twists [144, 290]. Two [322, 313, 198, 358, 47, 270, 352]. Type [138, 191, 17, 343, 174].

unbranched [57]. uniform [66, 262]. Unipotent [141, 311]. Uniqueness [187]. unit [293]. Unitary [140, 61]. Units [129, 288]. unity [336]. unlabelled [20]. untwisted [346]. Upper [69]. use [243, 220]. used [70]. uses [50]. Using [153, 83, 55, 287, 353, 2, 62, 35].

Václav [232]. Value [108]. valued [284]. values [298, 247]. vanishing [244]. variable [192]. variable-separated [192]. variation [344]. Variational [100, 156]. Varieties [95, 106, 218, 309]. vector [284]. vector-valued [284]. Verheul [102]. vertex [208]. via [279, 94, 170, 254]. Virtual [310]. Visualizing [354].

Waldspurger [238]. Weak [131, 241, 14]. Weierstrass [288]. Weight [132, 59, 73, 173]. weighted [245]. Weil [171, 47, 270, 40]. Wieferich [250]. windows [29]. without [364]. Witt [285].

yoga [179].

Zariski [214]. zero [343, 364, 168, 200]. zerofree [343]. Zeroes [331]. zeros [252, 40, 189]. Zeta [106, 229, 347, 43, 331, 227]. zetafunction [331]. ZF [62].

References

Conway:1998:TPG

[1] John H. Conway, Alexander Hulpke, and John McKay. On transitive permutation groups. LMS Journal of Computation and Mathematics, 1:1–8, 1998. CODEN ???? ISSN 1461-1570. URL http://

www.lms.ac.uk/jcm/1/lms1996-001/.

Leedham-Green:1998:SCU

[2] C. R. Leedham-Green and Leonard H. Soicher. Symbolic collection using Deep Thought. LMS Journal of Computation and Mathematics, 1:9–24, 1998. CODEN ???? ISSN 1461-1570. URL http://

www.lms.ac.uk/jcm/1/lms1997-002/.

Ellis:1998:CCH

[3] Graham Ellis. The computation of certain homotopical-functors. LMS Journal of Computation and Mathematics, 1:25–41, 1998. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/1/lms1997-004/.

Davies:1998:SEC

[4] E. B. Davies. Spectral enclosures and complex resonances for general selfadjoint operators. LMS Journal of Computation and Mathematics, 1:42–74, 1998. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/1/lms1997-005/.

Deriziotis:1998:CMF

[5] D. I. Deriziotis and C. P. Gotsis. The cuspidal modules of the finite general linear groups. LMS Journal of Computation and Mathematics, 1:75–108, 1998. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/1/lms1997-006/.

Sharp:1998:ARA

[6] Graham R. Sharp. Algorithmic recognition of actions of 2-homogeneous groups on pairs. LMS Journal of Computation and Mathematics, 1:109–147, 1998. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/1/lms1997-008/.

Russinoff:1998:MCP

[7] David M. Russinoff. A mechanically checked proof of IEEE compliance of the floating point multiplication, division and square root algorithms of the AMD-K7^{T M} processor. LMS Journal of Computation and Mathematics, 1:148–200, 1998. CODEN ???? ISSN 1461-1570.

URL http://www.lms.ac.uk/jcm/1/

lms1998-001/. Appendices A and B available to subscribers electronically (http://www.lms.ac.uk/jcm/1/lms98001/appendixa/ and http://www.lms.ac.uk/jcm/1/lms98001/appendixb/).

Sharp:1999:ARG

[8] Graham R. Sharp. Algorithmic recognition of group actions on orbitals. LMS Journal of Computation and Mathematics, 2:1–27, 1999. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/2/lms1998-006/.

Brown:1999:FCR

[9] Ronald Brown and Abdul Razak Salleh. Free crossed resolutions of groups and presentations of modules of identities among relations. LMS Journal of Computation and Mathematics, 2:28–61, 1999. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/2/lms1999-001/.

Cremona:1999:RBC

[10] J. E. Cremona. Reduction of binary cubic and quartic forms. LMS Journal of Computation and Mathematics, 2:62–92, 1999. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/2/lms1998-007/. See corrigen-

dum [34].

Ellis:1999:TDP

[11] Graham Ellis and Irina Kholodna. Three-dimensional presentations for the groups of order at most 30. LMS Journal of Computation and Mathematics, 2:93–117, 1999. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/2/lms1999-006/.

Galbraith:1999:CIB

[12] Steven D. Galbraith. Constructing isogenies between elliptic curves over finite fields. LMS Journal of Computation and Mathematics, 2:118–138, 1999. CODEN ???? ISSN 1461-1570. URL http://

                                        www.lms.ac.uk/jcm/2/lms1998-010/.

Davies:1999:CTS

[13] E. B. Davies. The computation of thresholds for Schrödinger operators. LMS Journal of Computation and Mathematics, 2:139–154, 1999. CODEN ???? ISSN 1461-1570. URL http://www.

                                        lms.ac.uk/jcm/2/lms1999-003/.

Adams:2000:FWN

[14] A. A. Adams. A formalisation of weak normalisation (with respect to permutations) of sequent calculus proofs. LMS Journal of Computation and Mathematics, 3:1–26, 2000. CODEN ???? ISSN

                                        1461-1570. URL http://www.lms.ac.
                                        uk/jcm/3/lms1999-009/.

Barclay:2000:TMD

[15] Graeme J. Barclay, David F. Griffiths, and Desmond J. Higham. Theta method dynamics. LMS Journal of Computation and Mathematics, 3:27–43, 2000. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/3/lms1999-004/.

Iserles:2000:DCG

[16] Arieh Iserles and Antonella Zanna. On the dimension of certain graded Lie algebras arising in geometric integration of differential equations. LMS Journal of Computation and Mathematics, 3:44–75, 2000. CODEN ???? ISSN

                                        1461-1570. URL http://www.lms.ac.
                                        uk/jcm/3/lms1999-011/.

Brown:2000:SCP

[17] B. M. Brown and M. S. P. Eastham. Spectral concentration for perturbed equations of harmonic oscillator type. LMS Journal of Computation and Mathematics, 3:76–85, 2000. CODEN ???? ISSN 1461-1570. URL http://

www.lms.ac.uk/jcm/3/lms1999-025/.

Borchert:2000:RAS

[18] Bernd Borchert, Lane A. Hemaspaandra, and Jörg Rothe. Restrictive acceptance suffices for equivalence problems. LMS Journal of Computation and Mathematics, 3:86–95, 2000. CODEN ???? ISSN 1461-1570. URL http://

www.lms.ac.uk/jcm/3/lms1999-012/.

Greenhill:2000:ARE

[19] Catherine Greenhill. An algorithm for recognising the exterior square of a multiset. LMS Journal of Computation and Mathematics, 3:96–116, 2000. CODEN ???? ISSN 1461-1570. URL http://

www.lms.ac.uk/jcm/3/lms1999-021/.

Goldberg:2000:CUS

[20] Leslie Ann Goldberg and Mark Jerrum. Counting unlabelled subtrees of a tree is #P-complete. LMS Journal of Computation and Mathematics, 3:117–124, 2000. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/3/lms1999-029/.

Einsiedler:2000:PSA

[21] Manfred Einsiedler, Graham Everest, and Thomas Ward. Primes in sequences associated to polynomials (after Lehmer). LMS Journal of Computation and Mathematics, 3:125–139, 2000. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/3/lms2000-004/.

Fleuriot:2000:MNR

[22] Jacques D. Fleuriot and Lawrence C. Paulson. Mechanizing nonstandard real analysis. LMS Journal of Computation and Mathematics, 3:140–190, 2000. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/3/lms1999-027/.

Bridges:2000:CAP

[23] Douglas Bridges and Robin Havea. A constructive analysis of a proof that the numerical range is convex. LMS Journal of Computation and Mathematics, 3: 191–206, 2000. CODEN ???? ISSN

                                        1461-1570. URL http://www.lms.ac.
                                        uk/jcm/3/lms1999-031/.

Kirk:2000:CAC

[24] N. P. Kirk. Computational aspects of classifying singularities. LMS Journal of Computation and Mathematics, 3:207– 228, 2000. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/3/lms1999-002/.

Howroyd:2000:DTA

[25] J. D. Howroyd. A domain-theoretic approach to integration in Hausdorff spaces. LMS Journal of Computation and Mathematics, 3:229–273, 2000. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/3/lms1999-010/.

Bleher:2000:SIG

[26] Frauke M. Bleher and Wolfgang Kimmerle. On the structure of integral group rings of sporadic groups. LMS Journal of Computation and Mathematics, 3: 274–306, 2000. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/3/lms2000-003/.

Foster:2000:EBH

[27] William H. Foster and Ilia Krasikov. Explicit bounds for Hermite polynomials in the oscillatory region. LMS Journal of Computation and Mathematics, 3: 307–314, 2000. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/3/lms2000-008/.

Baker:2000:NAE

[28] Christopher T. H. Baker and Evelyn Buckwar. Numerical analysis of explicit one-step methods for stochastic delay differential equations. LMS Journal of Computation and Mathematics, 3:315– 335, 2000. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/3/lms2000-002/.

Cavallar:2000:EW

[29] Stefania Cavallar and Franz Lemmermeyer. Euclidean windows. LMS Journal of Computation and Mathematics, 3: 336–355, 2000. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/3/lms2000-011/.

Einsiedler:2001:PED

[30] Manfred Einsiedler, Graham Everest, and Thomas Ward. Primes in elliptic divisibility sequences. LMS Journal of Computation and Mathematics, 4:1–13, 2001. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/4/lms2000-027/.

Draxler:2001:RDD

[31] Peter Dräxler. Representation-directed diamonds. LMS Journal of Computation and Mathematics, 4:13–20, 2001. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/4/lms2000-009/.

Hiss:2001:LDR

[32] Gerhard Hiss and Gunter Malle. Lowdimensional representations of quasisimple groups. LMS Journal of Computation and Mathematics, 4:22–63, 2001. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/4/lms2000-014/. See corrigenda

[46].

Detinko:2001:DFM

[33] A. Detinko. On deciding finiteness for matrix groups over fields of positive characteristic. LMS Journal of Computation and Mathematics, 4:64–72, 2001. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/4/lms2000-022/.

Cremona:2001:CRB

[34] J. E. Cremona. Corrigendum: “Reduction of binary cubic and quartic forms” [LMS J. Comput. Math. 2 (1999), 64– 94; MR 2000f:11040]. LMS Journal of Computation and Mathematics, 4:73, 2001. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/4/lms2001-012/. See [10].

Reinert:2001:CEU

[35] Birgit Reinert and Dirk Zeckzer. Coset enumeration using prefix Gröbner bases: an experimental approach. LMS Journal of Computation and Mathematics, 4:74–134, 2001. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/4/lms2000-010/.

Lubeck:2001:SDR

[36] Frank Lübeck. Small degree representations of finite Chevalley groups in defining characteristic. LMS Journal of Computation and Mathematics, 4:135–169, 2001. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/4/lms2000-015/.

Stein:2001:CMS

[37] William A. Stein and Helena A. Verrill. Cuspidal modular symbols are transportable. LMS Journal of Computation and Mathematics, 4:170–181, 2001. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/4/lms2001-006/.

Kluners:2001:DFE

[38] Jürgen Klüners and Gunter Malle. A database for field extensions of the rationals. LMS Journal of Computation and Mathematics, 4:182–196, 2001. CODEN ???? ISSN 1461-1570. URL http://

www.lms.ac.uk/jcm/4/lms2001-004/.

Chavarriga:2001:RD

[39] Javier Chavarriga, Jaume Llibre, and Jean Moulin Ollagnier. On a result of Darboux. LMS Journal of Computation and Mathematics, 4:197–210, 2001. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/4/lms2000-025/.

Omar:2002:NCL

[40] Sami Omar. Note on the contribution of low zeros to Weil’s explicit formula for minimal discriminants. LMS Journal of Computation and Mathematics, 5:1–6, 2002. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/5/lms2000-023/.

Simon:2002:CRE

[41] Denis Simon. Computing the rank of elliptic curves over number fields. LMS Journal of Computation and Mathematics, 5:7–17, 2002. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/5/lms2000-006/.

Muller:2002:CBT

[42] Jürgen Müller, Max Neunhöffer, Frank Röhr, and Robert Wilson. Completing the Brauer trees for the sporadic simple Lyons group. LMS Journal of Computation and Mathematics, 5:18–33, 2002. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/5/lms2001-013/.

Lauder:2002:CZF

[43] Alan G. B. Lauder and Daqing Wan. Computing zeta functions of Artin– Schreier curves over finite fields. LMS Journal of Computation and Mathematics, 5:34–55, 2002. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/5/lms2001-022/.

Gordon:2002:PCD

[44] Michael J. C. Gordon. Programming combinations of deduction and BDDbased symbolic calculation. LMS Journal of Computation and Mathematics, 5:56–76, 2002. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/5/lms2000-001/.

Breyer:2002:NMC

[45] L. A. Breyer and G. O. Roberts. A new method for coupling random fields. LMS Journal of Computation and Mathematics, 5:77–94, 2002. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/5/lms1999-007/.

Hiss:2002:CLD

[46] Gerhard Hiss and Gunter Malle. Corrigenda: Low-dimensional representations of quasi-simple groups. LMS Journal of Computation and Mathematics, 5:95– 126, 2002. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/5/lms2002-025/. See [32].

Maurer:2002:AGW

[47] Markus Maurer, Alfred Menezes, and Edlyn Treske. Analysis of the GHS Weil descent attack on the ECDLP over characteristic two finite fields of composite degree. LMS Journal of Computation and Mathematics, 5:127–174, 2002. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/5/lms2001-019/.

Wilson:2002:CCR

[48] Robert A. Wilson. Conjugacy class representatives in Fischer’s Baby Monster. LMS Journal of Computation and Mathematics, 5:175–180, 2002. CODEN ???? ISSN 1461-1570. URL http://www.

lms.ac.uk/jcm/5/lms2002-019/.

Harrison:2002:SIF

[49] K. Harrison, D. Page, and N. P. Smart. Software implementation of finite fields of characteristic three. LMS Journal of Computation and Mathematics, 5:181– 193, 2002. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/5/lms2002-002/.

Utting:2002:EGP

[50] Mark Utting, Peter Robinson, and Ray Nickson. Ergo 6: a generic proof engine that uses Prolog proof technology. LMS Journal of Computation and Mathematics, 5:194–219, 2002. CODEN ???? ISSN 1461-1570. URL http://www.

                                        lms.ac.uk/jcm/5/lms1999-032/.

Stoll:2002:MMC

[51] Michael Stoll and John E. Cremona. Minimal models for 2-coverings of elliptic curves. LMS Journal of Computation and Mathematics, 5:220–243, 2002. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/5/lms2002-013/.

Levitin:2003:RFT

[52] Michael Levitin and Rustem Yagudin. Range of the first three eigenvalues of the planar Dirichlet Laplacian. LMS Journal of Computation and Mathematics, 6:1–17, 2003. CODEN ???? ISSN

                                        1461-1570. URL http://www.lms.ac.
                                        uk/jcm/6/lms2002-005/.

Niesen:2003:PEG

[53] Jitse Niesen. A priori estimates for the global error committed by Runge–Kutta methods for a nonlinear oscillator. LMS Journal of Computation and Mathematics, 6:18–28, 2003. CODEN ???? ISSN

                                        1461-1570. URL http://www.lms.ac.
                                        uk/jcm/6/lms2001-024/.

Eick:2003:SPP

[54] B. Eick and B. Höfling. The solvable primitive permutation groups of degree at most 6560. LMS Journal of Computation and Mathematics, 6:29–39, 2003. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/6/lms2002-003/.

Stewart:2003:UPS

[55] Iain A. Stewart. Using program schemes to capture polynomial-time logically on certain classes of structures. LMS Journal of Computation and Mathematics, 6:40–67, 2003. CODEN ???? ISSN

1461-1570. URL http://www.lms.ac.
uk/jcm/6/lms2002-004/.

Bley:2003:NEC

[56] Werner Bley. Numerical evidence for a conjectural generalization of Hilbert’s Theorem 132. LMS Journal of Computation and Mathematics, 6:68–88, 2003. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/6/lms2002-021/. With an ap-

pendix by D. Kusnezow.

Kwak:2003:EBS

[57] Jin Ho Kwak, Jaeun Lee, and Alexander Mednykh. Enumerating branched surface coverings from unbranched ones. LMS Journal of Computation and Mathematics, 6:89–104, 2003. CODEN ???? ISSN 1461-1570. URL http://www.

lms.ac.uk/jcm/6/lms2001-016/.

deGraaf:2003:ACC

[58] Willem A. de Graaf. An algorithm to compute the canonical basis of an irreducible module over a quantized enveloping algebra. LMS Journal of Computation and Mathematics, 6:105–118, 2003. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/6/lms2002-032/.

An:2003:AWC

[59] Jianbei An, E. A. O’Brien, and R. A. Wilson. The Alperin weight conjecture and Dade’s conjecture for the simple group J_{4}. LMS Journal of Computation and Mathematics, 6:119–140, 2003. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/6/lms2003-003/.

Mao:2003:NSS

[60] Xuerong Mao. Numerical solutions of stochastic functional differential equations. LMS Journal of Computation and Mathematics, 6:141–161, 2003. CODEN ???? ISSN 1461-1570. URL http://

                                        www.lms.ac.uk/jcm/6/lms2002-027/.

Brooksbank:2003:FCR

[61] Peter A. Brooksbank. Fast constructive recognition of black-box unitary groups. LMS Journal of Computation and Mathematics, 6:162–197, 2003. CODEN ???? ISSN 1461-1570. URL http://www.

                                        lms.ac.uk/jcm/6/lms2003-008/.

Paulson:2003:RCA

[62] Lawrence C. Paulson. The relative consistency of the axiom of choice mechanized using Isabelle/ZF. LMS Journal of Computation and Mathematics, 6: 198–248, 2003. CODEN ???? ISSN

                                        1461-1570. URL http://www.lms.ac.
                                        uk/jcm/6/lms2003-001/.

Martin:2003:GBA

[63] Paul P. Martin and David Woodcock. Generalized blob algebras and alcove geometry. LMS Journal of Computation and Mathematics, 6:249–296, 2003. CODEN ???? ISSN 1461-

                                        1570. URL http://www.lms.ac.uk/
                                        jcm/6/lms2002-026/.

Higham:2003:EMS

[64] Desmond J. Higham, Xuerong Mao, and Andrew M. Stuart. Exponential meansquare stability of numerical solutions to stochastic differential equations. LMS Journal of Computation and Mathematics, 6:297–313, 2003. CODEN ???? ISSN 1461-1570. URL http://www.

lms.ac.uk/jcm/6/lms2003-014/.

Mossinghoff:2003:PRC

[65] Michael J. Mossinghoff. Polynomials with restricted coefficients and prescribed noncyclotomic factors. LMS Journal of Computation and Mathematics, 6:314–325, 2003. CODEN ???? ISSN 1461-1570. URL http://www.

lms.ac.uk/jcm/6/lms2003-013/.

Bridges:2003:SUC

[66] Douglas Bridges and Luminita Vita. Strong and uniform continuity — the uniform space case. LMS Journal of Computation and Mathematics, 6:326– 334, 2003. CODEN ???? ISSN 1461-

1570. URL http://www.lms.ac.uk/
jcm/6/lms2002-007/.

Davies:2004:CDS

[67] E. B. Davies. Computing the decay of a simple reversible sub-Markov semigroup. LMS Journal of Computation and Mathematics, 7:1–20, 2004. CODEN ???? ISSN 1461-1570.

McMurdy:2004:SMX

[68] Ken McMurdy. Stable model of X_{0}(125). LMS Journal of Computation and Mathematics, 7:21–36, 2004. CODEN ???? ISSN 1461-1570.

Cipu:2004:UBN

[69] Mihai Cipu. Upper bounds for norms of products of binomials. LMS Journal of Computation and Mathematics, 7:37–49, 2004. CODEN ???? ISSN 1461-1570.

Muzereau:2004:EBD

[70] A. Muzereau, N. P. Smart, and F. Vercauteren. The equivalence between the DHP and DLP for elliptic curves used in practical applications. LMS Journal of Computation and Mathematics, 7:50–72, 2004. CODEN ???? ISSN 1461-1570.

Neumann:2004:TFP

[71] Peter M. Neumann and Cheryl E. Praeger. On tensor-factorisation problems, I: the combinatorial problem. LMS Journal of Computation and Mathematics, 7:73–100, 2004. CODEN ???? ISSN 1461-1570.

Matthews:2004:ASB

[72] P. C. Matthews. Automating symmetrybreaking calculations. LMS Journal of Computation and Mathematics, 7:101– 119, 2004. CODEN ???? ISSN 1461-1570.

An:2004:AWC

[73] Jianbe An and R. A. Wilson. The Alperin weight conjecture and Uno’s conjecture for the Baby Monster B, p odd. LMS Journal of Computation and Mathematics, 7:120–166, 2004. CODEN ???? ISSN 1461-1570.

Hess:2004:GGA

[74] Florian Hess. Generalising the GHS attack on the elliptic curve discrete logarithm problem. LMS Journal of Computation and Mathematics, 7:167–192, 2004. CODEN ???? ISSN 1461-1570.

Le:2004:ERB

[75] Huiling Le. Estimation of Riemannian barycentres. LMS Journal of Computation and Mathematics, 7:193–200, 2004. CODEN ???? ISSN 1461-1570.

Galbraith:2004:EDD

[76] Steven D. Galbraith and Victor Rotger. Easy decision Diffie–Hellman groups. LMS Journal of Computation and Mathematics, 7:201–218, 2004. CODEN ???? ISSN 1461-1570.

Rowley:2004:CO

[77] Peter Rowley. Cross orbits. LMS Journal of Computation and Mathematics, 7: 219–265, 2004. CODEN ???? ISSN 1461-1570.

Campbell:2004:NEP

[78] Colin M. Campbell, George Havas, Colin Ramsay, and Edmund F. Robertson. Nice efficient presentations for all small simple groups and their covers. LMS Journal of Computation and Mathematics, 7:266–283, 2004. CODEN ???? ISSN 1461-1570.

Jacobs:2004:AAI

[79] Pascale Jacobs and Dimitri Leemans. An algorithmic analysis of the intersection property. LMS Journal of Computation and Mathematics, 7:284–299, 2004. CODEN ???? ISSN 1461-1570.

Vsemirnov:2004:HGI

[80] M. Vsemirnov. Hurwitz groups of intermediate rank. LMS Journal of Computation and Mathematics, 7:300–336, 2004. CODEN ???? ISSN 1461-1570.

Bates:2004:ICL

[81] Chris Bates and Peter Rowley. Involutions in Conway’s largest simple group.

LMS Journal of Computation and Mathematics, 7:337–351, 2004. CODEN ???? ISSN 1461-1570.

Bley:2005:PGR

[82] W. Bley and M. Endres. Picard groups and refined discrete logarithms. LMS Journal of Computation and Mathematics, 8:1–16, 2005. CODEN ???? ISSN 1461-1570.

Davies:2005:SBU

[83] E. B. Davies. Spectral bounds using higher-order numerical ranges. LMS Journal of Computation and Mathematics, 8:17–45, 2005. CODEN ???? ISSN 1461-1570.

Holt:2005:CMS

[84] Derek F. Holt and Colva M. Roney-Dougal. Constructing maximal subgroups of classical groups. LMS Journal of Computation and Mathematics, 8:46– 79, 2005. CODEN ???? ISSN 1461-1570.

Malle:2005:NFD

[85] Gunter Malle and David P. Roberts. Number fields with discriminant ±2^{a}3^{b} and Galois group A_{n} or S_{n}. LMS Journal of Computation and Mathematics, 8: 80–101, 2005. CODEN ???? ISSN 1461-1570.

Gutierrez:2005:HCE

[86] J. Gutierrez and T. Shaska. Hyperelliptic curves with extra involutions. LMS Journal of Computation and Mathematics, 8:102–115, 2005. CODEN ???? ISSN 1461-1570.

Chen:2005:DEA

[87] Imin Chen. A Diophantine equation associated to X_{0}(5). LMS Journal of Computation and Mathematics, 8:116–121, 2005. CODEN ???? ISSN 1461-1570.

Jansen:2005:MDF

[88] Christoph Jansen. The minimal degrees of faithful representations of the sporadic simple groups and their covering groups. LMS Journal of Computation and Mathematics, 8:122–144, 2005. CO-DEN ???? ISSN 1461-1570.

Howse:2005:SD

[89] John Howse, Gem Stapleton, and John Taylor. Spider diagrams. LMS Journal of Computation and Mathematics, 8: 145–194, 2005. CODEN ???? ISSN 1461-1570.

Charles:2005:CMP

[90] Denis Charles and Kristin Lauter. Computing modular polynomials. LMS Journal of Computation and Mathematics, 8: 195–204, 2005. CODEN ???? ISSN 1461-1570.

Barraclough:2005:CCR

[91] R. W. Barraclough and R. A. Wilson. Conjugacy class representatives in the Monster Group. LMS Journal of Computation and Mathematics, 8:205–216, 2005. CODEN ???? ISSN 1461-1570.

Ambrose:2005:GSB

[92] Sophie Ambrose, Max Neunhöffer, Cheryl E. Praeger, and Csaba Schneider. Generalised sifting in black-box groups. LMS Journal of Computation and Mathematics, 8:217–250, 2005. CODEN ???? ISSN 1461-1570.

Rodrigues:2005:NGA

[93] Rui C. Rodrigues, F. Silva Leite, and Janusz Jakubiak. A new geometric algorithm to generate smooth interpolating curves on Riemannian manifolds.

LMS Journal of Computation and Mathematics, 8:251–266, 2005. CODEN ???? ISSN 1461-1570.

Girard:2005:CSR

[94] Martine Girard and Leopoldo Kulesz. Computation of sets of rational points of genus-3 curves via the Dem’Janenko– Manin method. LMS Journal of Computation and Mathematics, 8:267–300, 2005. CODEN ???? ISSN 1461-1570.

Almeida:2005:HMI

[95] L. C. O. Almeida and S. C. Coutinho. On homogeneous minimal involutive varieties. LMS Journal of Computation and Mathematics, 8:301–315, 2005. CO-DEN ???? ISSN 1461-1570.

Goldberg:2006:IMB

[96] Leslie Ann Goldberg, Markus Jalsenius, Russell Martin, and Mike Paterson. Improved mixing bounds for the anti-ferromagnetic Potts model on Z^{2}. LMS Journal of Computation and Mathematics, 9:1–20, 2006. CODEN ???? ISSN 1461-1570.

Sinclair:2006:BNE

[97] Robert Sinclair and Minoru Tanaka. A bound on the number of endpoints of the cut locus. LMS Journal of Computation and Mathematics, 9:21–39, 2006. CO-DEN ???? ISSN 1461-1570.

Batty:2006:EPD

[98] Michael Batty, Andrew J. Duncan, and Samuel L. Braunstein. Extending the promise of the Deutsch–Jozsa–Høyer algorithm for finite groups. LMS Journal of Computation and Mathematics, 9:40– 63, 2006. CODEN ???? ISSN 1461-1570.

Granger:2006:SCA

[99] R. Granger, D. Page, and M. Stam. On small characteristic algebraic tori in pairing-based cryptography. LMS Journal of Computation and Mathematics, 9: 64–85, 2006. CODEN ???? ISSN 1461-1570.

Machado:2006:RMS

[100] Luís Machado, F. Silva Leite, and Knut Hüper. Riemannian means as solutions of variational problems. LMS Journal of Computation and Mathematics, 9:86– 103, 2006. CODEN ???? ISSN 1461-1570.

Detinko:2006:CNM

[101] A. S. Detinko and D. L. Flannery. Computing in nilpotent matrix groups. LMS Journal of Computation and Mathematics, 9:104–134, 2006. CODEN ???? ISSN 1461-1570.

Satoh:2006:PIR

[102] Takakazu Satoh. On polynomial interpolations related to Verheul homomorphisms. LMS Journal of Computation and Mathematics, 9:135–158, 2006. CO-DEN ???? ISSN 1461-1570.

Law:2006:RAL

[103] Maska Law, Alice C. Niemeyer, Cheryl E. Praeger, and Ákos Seress. A reduction algorithm for large-base primitive permutation groups. LMS Journal of Computation and Mathematics, 9:159–173, 2006. CODEN ???? ISSN 1461-1570.

Vaughan-Lee:2006:SLA

[104] Michael Vaughan-Lee. Simple Lie algebras of low dimension over GF(2).

LMS Journal of Computation and Mathematics, 9:174–192, 2006. CODEN ???? ISSN 1461-1570.

Roberts:2006:RMM

[105] A. J. Roberts. Resolving the multitude of microscale interactions accurately models stochastic partial differential equations. LMS Journal of Computation and Mathematics, 9:193–221, 2006. CODEN ???? ISSN 1461-1570.

Lauder:2006:RMC

[106] Alan G. B. Lauder. A recursive method for computing zeta functions of varieties. LMS Journal of Computation and Mathematics, 9:222–269, 2006. CODEN ???? ISSN 1461-1570.

Gornet:2006:LSI

[107] Ruth Gornet and Jeffrey McGowan. Lens spaces, isospectral on forms but not on functions. LMS Journal of Computation and Mathematics, 9:270–286, 2006. CODEN ???? ISSN 1461-1570.

Krupchyk:2006:SLC

[108] Katsiaryna Krupchyk and Jukka Tuomela. The Sharpiro–Lopatinskij condition for elliptic boundary value problems. LMS Journal of Computation and Mathematics, 9:287–329, 2006. CODEN ???? ISSN 1461-1570.

Jansons:2007:OCK

[109] Kalvis M. Jansons and Paul D. Metcalfe. Optimally coupling the Kolmogorov diffusion, and related optimal control problems. LMS Journal of Computation and Mathematics, 10:1–20, 2007. CODEN ???? ISSN 1461-1570.

Vincent:2007:NHC

[110] R. Vincent and A. E. Zalesski. Non-Hurwitz classical groups. LMS Journal of Computation and Mathematics, 10: 21–82, 2007. CODEN ???? ISSN 1461-1570.

Edalat:2007:DTA

[111] Abbas Edalat and Dirk Pattinson. A domain-theoretic account of Picard’s Theorem. LMS Journal of Computation and Mathematics, 10:83–118, 2007. CO-DEN ???? ISSN 1461-1570.

Nye:2007:CMI

[112] Tom M. W. Nye, Brad J. C. Baxter, and Walter R. Gilks. A covariance matrix inversion problem arising from the construction of phylogenetic trees. LMS Journal of Computation and Mathematics, 10:119–131, 2007. CODEN ???? ISSN 1461-1570.

Bradley:2007:IBS

[113] J. D. Bradley and P. E. Holmes. Improved bounds for the spread of sporadic groups. LMS Journal of Computation and Mathematics, 10:132–140, 2007. CODEN ???? ISSN 1461-1570.

Fraatz:2007:CIC

[114] Robert Fraatz. On the computation of integral closures of cyclic extensions of function fields. LMS Journal of Computation and Mathematics, 10:141–160, 2007. CODEN ???? ISSN 1461-1570.

Barraclough:2007:CTM

[115] R. W. Barraclough and R. A. Wilson.

The character table of a maximal subgroup of the Monster. LMS Journal of Computation and Mathematics, 10:161– 175, 2007. CODEN ???? ISSN 1461-1570.

McCusker:2007:CMS

[116] Guy McCusker. Categorical models of syntactic control of interference revisited, revisited. LMS Journal of Computation and Mathematics, 10:176–206, 2007. CODEN ???? ISSN 1461-1570.

Hubrechts:2007:PCF

[117] Hendrik Hubrechts. Point counting in families of hyperelliptic curves in characteristic 2. LMS Journal of Computation and Mathematics, 10:207–234, 2007. CODEN ???? ISSN 1461-1570.

Kloeden:2007:PCA

[118] P. E. Kloeden and A. Neuenkirch. The pathwise convergence of approximation schemes for stochastic differential equations. LMS Journal of Computation and Mathematics, 10:235–253, 2007. CO-DEN ???? ISSN 1461-1570.

Boissiere:2007:GSC

[119] Samuel Boissière and Marc A. Nieper-Wisskirchen. Generating series in the cohomology of Hilbert schemes of points on surfaces. LMS Journal of Computation and Mathematics, 10:254–270, 2007. CODEN ???? ISSN 1461-1570.

Fieker:2007:FIL

[120] Claus Fieker and Willem A. de Graaf. Finding integral linear dependencies of algebraic numbers and algebraic Lie algebras. LMS Journal of Computation and Mathematics, 10:271–287, 2007. CODEN ???? ISSN 1461-1570.

Davies:2007:ICD

[121] E. B. Davies. An indefinite convectiondiffusion operator. LMS Journal of Computation and Mathematics, 10:288– 306, 2007. CODEN ???? ISSN 1461-1570.

Salem:2007:FJG

[122] Fatima K. Abu Salem and Kamal khuri makdisi. Fast Jacobian group operations for C_{3,4} curves over a large finite field. LMS Journal of Computation and Mathematics, 10:307–328, 2007. CO-DEN ???? ISSN 1461-1570.

Ryba:2007:CSI

[123] A. J. E. Ryba. Construction of some irreducible subgroups of E_{8} and E_{6}. LMS Journal of Computation and Mathematics, 10:329–340, 2007. CODEN ???? ISSN 1461-1570.

Vaughan-Lee:2007:EG

[124] Michael Vaughan-Lee. On 4-Engel groups. LMS Journal of Computation and Mathematics, 10:341–353, 2007. CODEN ???? ISSN 1461-1570.

Bright:2007:BMO

[125] M. J. Bright, N. Bruin, E. V. Flynn, and A. Logan. The Brauer–Manin obstruction and III[2]. LMS Journal of Computation and Mathematics, 10:354–377, 2007. CODEN ???? ISSN 1461-1570.

Bosman:2007:PGG

[126] Johan Bosman. A polynomial with Galois groups SL_{2} (F_{16}). LMS Journal of Computation and Mathematics, 10:378– 388, 2007. CODEN ???? ISSN 1461-1570.

Lundow:2008:EAC

[127] Per Håkan Lundow and Klas Markström. Exact and approximate compression of transfer matrices for graph homomorphisms. LMS Journal of Computation and Mathematics, 11:1–14, 2008. CODEN ???? ISSN 1461-1570.

Muller:2008:ASS

[128] Jürgen Müller. On the action of the sporadic simple Baby Monster Group on its conjugacy class 2B. LMS Journal of Computation and Mathematics, 11:15– 27, 2008. CODEN ???? ISSN 1461-1570.

Bovdi:2008:TUI

[129] V. A. Bovdi, A. B. Konovalov, and S. Linton. Torsion units in integral group ring of the Mathieu simple group M_{22}. LMS Journal of Computation and Mathematics, 11:28–39, 2008. CODEN ???? ISSN 1461-1570.

Harvey:2008:ECA

[130] David Harvey. Efficient computation of p-adic heights. LMS Journal of Computation and Mathematics, 11:40–59, 2008. CODEN ???? ISSN 1461-1570.

Buckwar:2008:WCE

[131] Evelyn Buckwar, Rachel Kuske, Salah-Eldin Mohammed, and Tony Shardlow. Weak convergence of the Euler scheme for stochastic differential delay equations. LMS Journal of Computation and Mathematics, 11:60–99, 2008. CODEN ???? ISSN 1461-1570.

An:2008:AWC

[132] Jianbei An, John J. Cannon, E. A. O’Brien, and W. R. Unger. The Alperin weight conjecture and Dade’s Conjecture for the simple group Fi′_{24}. LMS Journal of Computation and Mathematics, 11:100–145, 2008. CODEN ???? ISSN 1461-1570.

McKay:2008:DRF

[133] J. McKay and David Sevilla. Decomposing replicable functions. LMS Journal of Computation and Mathematics, 11:146– 171, 2008. CODEN ???? ISSN 1461-1570.

Lauder:2008:REC

[134] Alan G. B. Lauder. Ranks of elliptic curves over function fields. LMS Journal of Computation and Mathematics, 11:172–212, 2008. CODEN ???? ISSN 1461-1570.

Noeske:2008:MSM

[135] Felix Noeske. Matching simple modules of condensed algebras. LMS Journal of Computation and Mathematics, 11:213– 222, 2008. CODEN ???? ISSN 1461-1570.

Holt:2008:CCS

[136] Derek F. Holt and Mark J. Stather.

Computing a chief series and the soluble radical of a matrix group over a finite field. LMS Journal of Computation and Mathematics, 11:223–251, 2008. CO-DEN ???? ISSN 1461-1570.

Neunhoffer:2008:CMP

[137] Max Neunhöffer and Cheryl E. Praeger.

Computing minimal polynomials of matrices. LMS Journal of Computation and Mathematics, 11:252–279, 2008. CO-DEN ???? ISSN 1461-1570.

deGraaf:2008:CNO

[138] Willem A. de Graaf. Computing with nilpotent orbits in simple Lie algebras of exceptional type. LMS Journal of Computation and Mathematics, 11:280–297, 2008. CODEN ???? ISSN 1461-1570.

Mahe:2008:HSP

[139] Valéry Mahé. Hilbert’s Seventeenth Problem and hyperelliptic curves. LMS Journal of Computation and Mathematics, 11:298–325, 2008. CODEN ???? ISSN 1461-1570.

Loeffler:2008:ECA

[140] David Loeffler. Explicit calculations of automorphic forms for definite unitary groups. LMS Journal of Computation and Mathematics, 11:326–342, 2008. CODEN ???? ISSN 1461-1570.

Cohen:2008:CUR

[141] Arjeh M. Cohen, Sergei Haller, and Scott H. Murray. Computing in unipotent and reductive algebraic groups. LMS Journal of Computation and Mathematics, 11:343–366, 2008. CODEN ???? ISSN 1461-1570.

Stoll:2008:RCU

[142] Michael Stoll. Rational 6-cycles under iteration of quadratic polynomials. LMS Journal of Computation and Mathematics, 11:367–380, 2008. CODEN ???? ISSN 1461-1570.

Himstedt:2009:CTB

[143] Frank Himstedt and Shih-Chang Huang. Character table of a Borel subgroup of the Ree groups ^{2}F_{4}(q^{2}). LMS Journal of Computation and Mathematics, 12:1–53, 2009. CODEN ???? ISSN 1461-1570.

Everest:2009:PDT

[144] Graham Everest, Patrick Ingram, and Shaun Stevens. Primitive divisors on twists of Fermat’s Cubic. LMS Journal of Computation and Mathematics, 12:54–81, 2009. CODEN ???? ISSN 1461-1570.

Babai:2009:NRE

[145] László Babai, Péter P. Pálfy, and Jan Saxl. On the number of p-regular elements in finite simple groups. LMS Journal of Computation and Mathematics, 12:82–119, 2009. CODEN ???? ISSN 1461-1570.

Rowley:2009:CGM

[146] Peter Rowley. The Chamber graph of the M_{24} maximal 2-local geometry. LMS Journal of Computation and Mathematics, 12:120–143, 2009. CODEN ???? ISSN 1461-1570.

Varilly-Alvarado:2009:ALM

[147] Anthony Várilly-Alvarado and David Zywina. Arithmetic E_{8} lattices with maximal Galois action. LMS Journal of Computation and Mathematics, 12:144– 165, 2009. CODEN ???? ISSN 1461-1570.

Bley:2009:CRA

[148] Werner Bley and Stephen M. J. Wilson. Computations in relative algebraic K-groups. LMS Journal of Computation and Mathematics, 12:166–194, 2009. CODEN ???? ISSN 1461-1570.

Jalsenius:2009:SSM

[149] Markus Jalsenius. Strong spatial mixing and rapid mixing with five colours for the Kagome lattice. LMS Journal of Computation and Mathematics, 12:195– 227, 2009. CODEN ???? ISSN 1461-1570.

Huang:2009:GDS

[150] Ming-Deh Huang and Wayne Raskind.

Global duality, signature calculus and the discrete logarithm problem. LMS Journal of Computation and Mathematics, 12:228–263, 2009. CODEN ???? ISSN 1461-1570.

Cummins:2009:CCC

[151] C. J. Cummins. On conjugacy classes of congruence subgroups of PSL(2, R). LMS Journal of Computation and Mathematics, 12:264–274, 2009. CODEN ???? ISSN 1461-1570.

Katz:2009:TI

[152] Eric Katz, Hannah Markwig, and Thomas Markwig. The tropical jinvariant. LMS Journal of Computation and Mathematics, 12:275–294, 2009. CODEN ???? ISSN 1461-1570.

Chatel:2009:PCA

[153] Gweltaz Chatel and David Lubicz. A point counting algorithm using cohomology with compact support. LMS Journal of Computation and Mathematics, 12:295–325, 2009. CODEN ???? ISSN 1461-1570.

Broker:2009:MPG

[154] Reinier Bröker and Kristin Lauter. Modular polynomials for genus 2. LMS Journal of Computation and Mathematics, 12:326–339, 2009. CODEN ???? ISSN 1461-1570.

Smith:2010:CIS

[155] Kevin Smith. On complete interpolating sequences and sampling expansions. LMS Journal of Computation and Mathematics, 13:1–9, 2010. CODEN ???? ISSN 1461-1570.

Boulton:2010:NVC

[156] Lyonell Boulton and Nabile Boussaid. Non-variational computation of the eigenstates of Dirac operators with radially symmetric potentials. LMS Journal of Computation and Mathematics, 13:10–32, 2010. CODEN ???? ISSN 1461-1570.

Kasprzyk:2010:CCT

[157] Alexander M. Kasprzyk, Maximilian Kreuzer, and Benjamin Nill. On the combinatorial classification of toric log del Pezzo surfaces. LMS Journal of Computation and Mathematics, 13:33– 46, 2010. CODEN ???? ISSN 1461-1570.

Muller:2010:EKS

[158] Jan Steffen Müller. Explicit Kummer surface formulas for arbitrary characteristic. LMS Journal of Computation and Mathematics, 13:47–64, 2010. CODEN ???? ISSN 1461-1570.

Brown:2010:EEE

[159] B. Malcolm Brown, Matthias Langer, Marco Marletta, Christiane Tretter, and Markus Wagenhofer. Eigenvalue enclosures and exclosures for non-self-adjoint problems in hydrodynamics. LMS Journal of Computation and Mathematics, 13:65–81, 2010. CODEN ???? ISSN 1461-1570.

Barraclough:2010:CTG

[160] R. W. Barraclough. The character table of a group of shape (2 × 2 · G): 2. LMS Journal of Computation and Mathematics, 13:82–89, 2010. CODEN ???? ISSN 1461-1570.

Himstedt:2010:CTM

[161] Frank Himstedt and Shih-Chang Huang.

Character tables of the maximal parabolic subgroups of the Ree groups 2F_{4}(q^{2}). LMS Journal of Computation and Mathematics, 13:90–110, 2010. CO-DEN ???? ISSN 1461-1570.

Perry:2010:EBC

[162] John Perry. An extension of Buchberger’s criteria for Gröbner basis decision. LMS Journal of Computation and Mathematics, 13:111–129, 2010. CO-DEN ???? ISSN 1461-1570.

Hare:2010:LBG

[163] Kevin G. Hare and Nikita Sidorov. A lower bound for Garsia’s entropy for certain Bernoulli convolutions. LMS Journal of Computation and Mathematics, 13:130–143, 2010. CODEN ???? ISSN 1461-1570.

Kontogeorgis:2010:AHM

[164] Aristides Kontogeorgis and Yifan Yang. Automorphisms of hyperelliptic modular curves X_{0}(N) in positive characteristic. LMS Journal of Computation and Mathematics, 13:144–163, 2010. CO-DEN ???? ISSN 1461-1570.

Holt:2010:CMS

[165] Derek F. Holt and Colva M. Roney-Dougal. Constructing maximal subgroups of orthogonal groups. LMS Journal of Computation and Mathematics, 13:164–191, 2010. CODEN ???? ISSN 1461-1570.

Ritzenthaler:2010:ECS

[166] Christophe Ritzenthaler. Explicit computations of Serre’s obstruction for genus-3 curves and application to optimal curves. LMS Journal of Computation and Mathematics, 13:192–207, 2010. CODEN ???? ISSN 1461-1570.

Ninomiya:2010:AKA

[167] Mariko Ninomiya. Application of the Kusuoka approximation with a treebased branching algorithm to the pricing of interest-rate derivatives under the HJM model. LMS Journal of Computation and Mathematics, 13:208–221, 2010. CODEN ???? ISSN 1461-1570.

Cummins:2010:FDG

[168] C. J. Cummins. Fundamental domains for genus-zero and genus-one congruence subgroups. LMS Journal of Computation and Mathematics, 13:222–245, 2010. CODEN ???? ISSN 1461-1570.

Berthelsen:2010:PPS

[169] Kasper K. Berthelsen, Laird A. Breyer, and Gareth O. Roberts. Perfect posterior simulation for mixture and hidden Markov models. LMS Journal of Computation and Mathematics, 13:246–259, 2010. CODEN ???? ISSN 1461-1570.

Hartung:2010:ASM

[170] René Hartung. Approximating the Schur multiplier of certain infinitely presented groups via nilpotent quotients. LMS Journal of Computation and Mathematics, 13:260–271, 2010. CODEN ???? ISSN 1461-1570.

Bruin:2010:MWS

[171] Nils Bruin and Michael Stoll. The Mordell–Weil sieve: proving nonexistence of rational points on curves. LMS Journal of Computation and Mathematics, 13:272–306, 2010. CODEN ???? ISSN 1461-1570.

Rowley:2010:NSB

[172] Peter Rowley and Paul Taylor. Normalizers of 2-subgroups in black-box groups. LMS Journal of Computation and Mathematics, 13:307–319, 2010. CODEN ???? ISSN 1461-1570.

An:2010:AWC

[173] Jianbei An and R. A. Wilson. The Alperin weight conjecture and Uno’s conjecture for the Monster M, p odd. LMS Journal of Computation and Mathematics, 13:320–356, 2010. CODEN ???? ISSN 1461-1570.

Goodwin:2010:DRF

[174] Simon M. Goodwin, Gerhard Röhrle, and Glenn Ubly. On 1-dimensional representations of finite W -algebras associated to simple Lie algebras of exceptional type. LMS Journal of Computation and Mathematics, 13:357–369, 2010. CODEN ???? ISSN 1461-1570.

Jain:2010:PLH

[175] Sonal Jain. Points of low height on elliptic surfaces with torsion. LMS Journal of Computation and Mathematics, 13:370– 387, 2010. CODEN ???? ISSN 1461-1570.

Baber:2010:MDT

[176] Rahil Baber, J. Robert Johnson, and John Talbot. The minimal density of triangles in tripartite graphs. LMS Journal of Computation and Mathematics, 13:388–413, 2010. CODEN ???? ISSN 1461-1570.

Calude:2010:CFC

[177] Cristian S. Calude and Elena Calude. The complexity of the four colour theorem. LMS Journal of Computation and Mathematics, 13:414–425, 2010. CO-DEN ???? ISSN 1461-1570.

Malle:2010:CRH

[178] Gunter Malle and Jean Michel. Constructing representations of Hecke algebras for complex reflection groups. LMS Journal of Computation and Mathematics, 13:426–450, 2010. CODEN ???? ISSN 1461-1570.

Fisher:2010:YCT

[179] Tom Fisher, Edward F. Schaefer, and Michael Stoll. The yoga of the Cassels– Tate pairing. LMS Journal of Computation and Mathematics, 13:451–460, 2010. CODEN ???? ISSN 1461-1570.

Yang:2010:RME

[180] Le Yang. Riemannian median and its estimation. LMS Journal of Computation and Mathematics, 13:461–479, 2010. CODEN ???? ISSN 1461-1570.

Lauder:2011:DLF

[181] Alan G. B. Lauder. Degenerations and limit Frobenius structures in rigid cohomology. LMS Journal of Computation and Mathematics, 14:1–33, 2011. CO-DEN ???? ISSN 1461-1570.

Freden:2011:GBS

[182] Eric M. Freden, Teresa Knudson, and Jennifer Schofield. Growth in Baumslag–Solitar groups I: subgroups and rationality. LMS Journal of Computation and Mathematics, 14:34–71, 2011. CODEN ???? ISSN 1461-1570.

Taylor:2011:DBS

[183] Alan Taylor, J. Keith Vass, and Desmond J. Higham. Discovering bipartite substructure in directed networks. LMS Journal of Computation and Mathematics, 14:72–86, 2011. CODEN ???? ISSN 1461-1570.

Rossmann:2011:PTF

[184] Tobias Rossmann. Primitivity testing of finite nilpotent linear groups. LMS Journal of Computation and Mathematics, 14:87–98, 2011. CODEN ???? ISSN 1461-1570.

Mihoubi:2011:MCF

[185] Douadi Mihoubi. Muller condition and fairness on multitransition systems. LMS Journal of Computation and Mathematics, 14:99–107, 2011. CODEN ???? ISSN 1461-1570.

Broker:2011:ACI

[186] Reinier Bröker. p-adic class invariants.

LMS Journal of Computation and Mathematics, 14:108–126, 2011. CODEN ???? ISSN 1461-1570.

Diener:2011:UCE

[187] H. Diener and P. Schuster. Uniqueness, continuity and the existence of implicit functions in constructive analysis. LMS Journal of Computation and Mathematics, 14:127–136, 2011. CODEN ???? ISSN 1461-1570.

Smith:2011:CSC

[188] Kevin Smith. On the completeness of sets of complex exponentials. LMS Journal of Computation and Mathematics, 14:137–139, 2011. CODEN ???? ISSN 1461-1570.

Omar:2011:ZDF

[189] Sami Omar, Raouf Ouni, and Kamel Mazhouda. On the zeros of Dirichlet L-functions. LMS Journal of Computation and Mathematics, 14:140–154, 2011. CODEN ???? ISSN 1461-1570.

Sabeti:2011:NSE

[190] Rostam Sabeti. Numerical-symbolic exact irreducible decomposition of cyclic-12. LMS Journal of Computation and Mathematics, 14:155–172, 2011. CO-DEN ???? ISSN 1461-1570.

Breuer:2011:CCT

[191] Thomas Breuer. Computing character tables of groups of type M. G. A. LMS Journal of Computation and Mathematics, 14:173–178, 2011. CODEN ???? ISSN 1461-1570.

Smith:2011:FEI

[192] Benjamin Smith. Families of explicitly isogenous Jacobians of variableseparated curves. LMS Journal of Computation and Mathematics, 14:179–199, 2011. CODEN ???? ISSN 1461-1570.

Johnson:2011:SSS

[193] Kenneth W. Johnson, Michael K.

Kinyon, Gábor P. Nagy, and Petr Vojtechovský. Searching for small simple automorphic loops. LMS Journal of Computation and Mathematics, 14:200–213, 2011. CODEN ???? ISSN 1461-1570.

Lauder:2011:CCA

[194] Alan G. B. Lauder. Computations with classical and p-adic modular forms. LMS Journal of Computation and Mathematics, 14:214–231, 2011. CODEN ???? ISSN 1461-1570.

Neumann:2011:NTP

[195] Peter M. Neumann. A note on the triple product property for subsets of finite groups. LMS Journal of Computation and Mathematics, 14:232–237, 2011. CODEN ???? ISSN 1461-1570.

Rowley:2011:IJS

[196] Peter Rowley and Paul Taylor. Involutions in Janko’s simple group J_{4}. LMS Journal of Computation and Mathematics, 14:238–253, 2011. CODEN ???? ISSN 1461-1570.

Park:2011:ENF

[197] Jun H. Park, Boris Rozovskii, and Richard B. Sowers. Efficient nonlinear filtering of a singularly perturbed stochastic hybrid system. LMS Journal of Computation and Mathematics, 14:254–270, 2011. CODEN ???? ISSN 1461-1570.

Chlouveraki:2011:DMD

[198] Maria Chlouveraki and Hyohe Miyachi.

Decomposition matrices for d-Harish-Chandra series: the exceptional rank two cases. LMS Journal of Computation and Mathematics, 14:271–290, 2011. CODEN ???? ISSN 1461-1570.

Eilbeck:2011:AFA

[199] J. C. Eilbeck, M. England, and Y. Ônishi. Abelian functions associated with genus three algebraic curves. LMS Journal of Computation and Mathematics, 14:291–326, 2011. CODEN ???? ISSN 1461-1570.

Miller:2011:PBS

[200] Robert L. Miller. Proving the Birch and Swinnerton-Dyer conjecture for specific elliptic curves of analytic rank zero and one. LMS Journal of Computation and Mathematics, 14:327–350, 2011. CO-DEN ???? ISSN 1461-1570.

Moussa:2012:DRH

[201] Abdelaziz Aït Moussa and Loubna Zlaïji. Dimension reduction and homogenization of random degenerate operators. Part I. LMS Journal of Computation and Mathematics, 15:1–22, 2012. CODEN ???? ISSN 1461-1570.

Arnaudon:2012:MMF

[202] Marc Arnaudon and Frank Nielsen. Medians and means in Finsler geometry. LMS Journal of Computation and Mathematics, 15:23–37, 2012. CODEN ???? ISSN 1461-1570.

Pak:2012:TCG

[203] Igor Pak. Testing commutativity of a group and the power of randomization. LMS Journal of Computation and Mathematics, 15:38–43, 2012. CODEN ???? ISSN 1461-1570.

Dunand:2012:MIC

[204] Clément Dunand. On modular inverses of cyclotomic polynomials and the magnitude of their coefficients. LMS Journal of Computation and Mathematics, 15:44–58, 2012. CODEN ???? ISSN 1461-1570.

Islam:2012:PTP

[205] Asim Islam. Products of three pairwise coprime integers in short intervals. LMS Journal of Computation and Mathematics, 15:59–70, 2012. CODEN ???? ISSN 1461-1570.

Berkolaiko:2012:ASA

[206] Gregory Berkolaiko, Evelyn Buckwar, Cónall Kelly, and Alexandra Rodkina. Almost sure asymptotic stability analysis of the θ-Maruyama method applied to a test system with stabilising and destabilising stochastic perturbations. LMS Journal of Computation and Mathematics, 15:71–83, 2012. CODEN ???? ISSN 1461-1570. See corrigendum [243].

Elsenhans:2012:KSC

[207] Andreas-Stephan Elsenhans and Jörg Jahnel. Kummer surfaces and the computation of the Picard group. LMS Journal of Computation and Mathematics, 15:84–100, 2012. CODEN ???? ISSN 1461-1570.

Carbone:2012:GAS

[208] Lisa Carbone, Riikka Kangaslampi, and Alina Vdovina. Groups acting simply transitively on vertex sets of hyperbolic triangular buildings. LMS Journal of Computation and Mathematics, 15:101– 112, 2012. CODEN ???? ISSN 1461-1570.

Kilford:2012:SOA

[209] L. J. P. Kilford and Ken McMurdy.

Slopes of the U_{7} operator acting on a space of overconvergent modular forms. LMS Journal of Computation and Mathematics, 15:113–139, 2012. CODEN ???? ISSN 1461-1570.

Lux:2012:CHS

[210] Klaus Lux, Max Neunhöffer, and Felix Noeske. Condensation of homomorphism spaces. LMS Journal of Computation and Mathematics, 15:140–157, 2012. CODEN ???? ISSN 1461-1570.

Chen:2012:EBC

[211] Imin Chen. On the equations a^{2} − 2b^{6} = c^{p} and a^{2} − 2 = c^{p}. LMS Journal of Computation and Mathematics, 15:158– 171, 2012. CODEN ???? ISSN 1461-1570.

Sutherland:2012:ACM

[212] Andrew V. Sutherland. Accelerating the CM method. LMS Journal of Computation and Mathematics, 15:172–204, 2012. CODEN ???? ISSN 1461-1570.

Eick:2012:CHS

[213] Bettina Eick and Alexander Hulpke. Computing Hall subgroups of finite groups. LMS Journal of Computation and Mathematics, 15:205–218, 2012. CODEN ???? ISSN 1461-1570.

Bauer:2012:ZCS

[214] Thomas Bauer and David Schmitz. Zariski chambers on surfaces of high Picard number. LMS Journal of Computation and Mathematics, 15:219–230, 2012. CODEN ???? ISSN 1461-1570.

Geck:2012:PCF

[215] Meinolf Geck. PyCox: computing with (finite) Coxeter groups and Iwahori– Hecke algebras. LMS Journal of Computation and Mathematics, 15:231–256, 2012. CODEN ???? ISSN 1461-1570.

Hiss:2012:BCS

[216] Gerhard Hiss, Jürgen Müller, Felix Noeske, and Jon Thackray. The Brauer characters of the sporadic simple Harada–Norton group and its automorphism group in characteristics 2 and 3. LMS Journal of Computation and Mathematics, 15:257–280, 2012. CO-DEN ???? ISSN 1461-1570.

Morgan:2012:GGC

[217] Kerri Morgan. Galois groups of chromatic polynomials. LMS Journal of Computation and Mathematics, 15:281– 307, 2012. CODEN ???? ISSN 1461-1570.

Arene:2012:CAL

[218] Christophe Arene, David Kohel, and Christophe Ritzenthaler. Complete addition laws on abelian varieties. LMS Journal of Computation and Mathematics, 15:308–316, 2012. CODEN ???? ISSN 1461-1570.

Sutherland:2012:ISE

[219] Andrew V. Sutherland. Identifying supersingular elliptic curves. LMS Journal of Computation and Mathematics, 15:317–325, 2012. CODEN ???? ISSN 1461-1570.

Diem:2012:UES

[220] Claus Diem. On the use of expansion series for stream ciphers. LMS Journal of Computation and Mathematics, 15:326– 340, 2012. CODEN ???? ISSN 1461-1570.

Johansson:2012:EIH

[221] Fredrik Johansson. Efficient implementation of the Hardy–Ramanujan– Rademacher formula. LMS Journal of Computation and Mathematics, 15:341– 359, 2012. CODEN ???? ISSN 1461-1570.

Siurys:2012:LRS

[222] Jonas Siurys. A linear recurrence sequence of composite numbers. LMS Journal of Computation and Mathematics, 15:360–373, 2012. CODEN ???? ISSN 1461-1570.

Noeske:2012:MSM

[223] Felix Noeske. Matching simple modules of condensation algebras. LMS Journal of Computation and Mathematics, 15: 374–384, 2012. CODEN ???? ISSN 1461-1570.

Magaard:2012:ACC

[224] K. Magaard and R. A. Wilson. Algorithmic construction of Chevalley bases. LMS Journal of Computation and Mathematics, 15:436–443, 2012. CODEN ???? ISSN 1461-1570.

Bettayeb:2012:ELM

[225] Abdellatif Bettayeb. Examples of linear multi-box splines. LMS Journal of Computation and Mathematics, 15:444–462, 2012. CODEN ???? ISSN 1461-1570.

Pakdemirli:2013:CPT

[226] M. Pakdemirli and G. Sari. A comprehensive perturbation theorem for estimating magnitudes of roots of polynomials. LMS Journal of Computation and Mathematics, 16:1–8, 2013. CODEN ???? ISSN 1461-1570.

Sperber:2013:CZF

[227] Steven Sperber and John Voight. Computing zeta functions of nondegenerate hypersurfaces with few monomials. LMS Journal of Computation and Mathematics, 16:9–44, 2013. CODEN ???? ISSN 1461-1570.

Calvi:2013:CML

[228] J.-P. Calvi and V. M. Phung. On the continuity of multivariate Lagrange interpolation at natural lattices. LMS Journal of Computation and Mathematics, 16:45–60, 2013. CODEN ???? ISSN 1461-1570.

Forster:2013:CBS

[229] B. Forster, R. Garunkstis, P. Massopust, and J. Steuding. Complex B-splines and Hurwitz zeta functions. LMS Journal of Computation and Mathematics, 16:61– 77, 2013. CODEN ???? ISSN 1461-1570.

Booker:2013:BAB

[230] Andrew R. Booker, Andreas Strömbergsson, and Holger Then. Bounds and algorithms for the K-Bessel function of imaginary order. LMS Journal of Computation and Mathematics, 16:78–108, 2013. CODEN ???? ISSN 1461-1570.

Brown:2013:SNC

[231] Gavin Brown and Alexander M.

Kasprzyk. Seven new champion linear codes. LMS Journal of Computation and Mathematics, 16:109–117, 2013. CO-DEN ???? ISSN 1461-1570.

Lemmermeyer:2013:VSQ

[232] F. Lemmermeyer. Václav Simerka:

quadratic forms and factorization. LMS Journal of Computation and Mathematics, 16:118–129, 2013. CODEN ???? ISSN 1461-1570.

Jones:2013:MSN

[233] John W. Jones. Minimal solvable nonic fields. LMS Journal of Computation and Mathematics, 16:130–138, 2013. CO-DEN ???? ISSN 1461-1570.

Bauch:2013:COF

[234] Jens-Dietrich Bauch, Enric Nart, and Hayden D. Stainsby. Complexity of OM factorizations of polynomials over local fields. LMS Journal of Computation and Mathematics, 16:139–171, 2013. CO-DEN ???? ISSN 1461-1570.

Lorch:2013:SCG

[235] David Lorch and Markus Kirschmer. Single-class genera of positive integral lattices. LMS Journal of Computation and Mathematics, 16:172–186, 2013. CODEN ???? ISSN 1461-1570.

Rahm:2013:LOC

[236] Alexander D. Rahm and Mehmet Haluk Sengün. On level one cuspidal Bianchi modular forms. LMS Journal of Computation and Mathematics, 16:187–199, 2013. CODEN ???? ISSN 1461-1570.

Gehre:2013:CBP

[237] Dominic Gehre, Judith Kreuzer, and Martin Raum. Computing Borcherds products. LMS Journal of Computation and Mathematics, 16:200–215, 2013. CODEN ???? ISSN 1461-1570.

Purkait:2013:EAW

[238] Soma Purkait. Explicit application of Waldspurger’s theorem. LMS Journal of Computation and Mathematics, 16:216– 245, 2013. CODEN ???? ISSN 1461-1570.

Citro:2013:CLO

[239] Craig Citro and Alexandru Ghitza. Computing level one Hecke eigensystems (mod p). LMS Journal of Computation and Mathematics, 16:246–270, 2013. CODEN ???? ISSN 1461-1570.

He:2013:MSD

[240] Yang-Hui He, John McKay, and James Read. Modular subgroups, dessins d’enfants and elliptic K3 surfaces. LMS Journal of Computation and Mathematics, 16:271–318, 2013. CODEN ???? ISSN 1461-1570.

Zhang:2013:WAS

[241] Hua Zhang. Weak approximation of stochastic differential delay equations for bounded measurable function. LMS Journal of Computation and Mathematics, 16:319–343, 2013. CODEN ???? ISSN 1461-1570.

Rahm:2013:HTA

[242] Alexander D. Rahm. Higher torsion in the Abelianization of the full Bianchi groups. LMS Journal of Computation and Mathematics, 16:344–365, 2013. CODEN ???? ISSN 1461-1570.

Berkolaiko:2013:CUD

[243] Gregory Berkolaiko, Evelyn Buckwar, Cónall Kelly, and Alexandra Rodkina. Corrigendum: On the use of a discrete form of the Itô formula in the article ‘Almost sure asymptotic stability analysis of the θ-Maruyama method applied to a test system with stabilising and destabilising stochastic perturbations’. LMS Journal of Computation and Mathematics, 16:366–372, 2013. CODEN ???? ISSN 1461-1570. See [206].

Dumnicki:2013:VTS

[244] M. Dumnicki, T. Szemberg, and H. Tutaj-Gasińska. A vanishing theorem and symbolic powers of planar point ideals. LMS Journal of Computation and Mathematics, 16:373–387, 2013. CO-DEN ???? ISSN 1461-1570.

Izgi:2013:ACC

[245] Aydin Izgi. Approximation by a composition of Chlodowsky operators and Százs–Durrmeyer operators on weighted spaces. LMS Journal of Computation and Mathematics, 16:388–397, 2013. CODEN ???? ISSN 1461-1570.

Dabbaghian:2013:CCG

[246] Vahid Dabbaghian and John D. Dixon. Computing characters of groups with central subgroups. LMS Journal of Computation and Mathematics, 16:398– 406, 2013. CODEN ???? ISSN 1461-1570.

Enge:2013:SVM

[247] Andreas Enge and Reinhard Schertz. Singular values of multiple eta-quotients for ramified primes. LMS Journal of Computation and Mathematics, 16:407– 418, 2013. CODEN ???? ISSN 1461-1570.

Jambor:2013:MGS

[248] Sebastian Jambor. The minimal generating sets of PSL(2, p) of size four. LMS Journal of Computation and Mathematics, 16:419–423, 2013. CODEN ???? ISSN 1461-1570.

Franc:2014:CFD

[249] Cameron Franc and Marc Masdeu. Computing fundamental domains for the Bruhat–Tits tree for GL_{2}(Q_{p}), p-adic automorphic forms, and the canonical embedding of Shimura curves. LMS Journal of Computation and Mathematics, 17(1):1–23, 2014. CODEN ???? ISSN 1461-1570.

Borwein:2014:WPB

[250] Peter Borwein and Michael J. Mossinghoff. Wieferich pairs and Barker sequences, II. LMS Journal of Computation and Mathematics, 17(1):24–32, 2014. CODEN ???? ISSN 1461-1570.

Wilson:2014:CSI

[251] Robert A. Wilson. Classification of subgroups isomorphic to PSL_{2}(27) in the Monster. LMS Journal of Computation and Mathematics, 17(1):33–46, 2014. CODEN ???? ISSN 1461-1570.

Jooste:2014:BZM

[252] A. Jooste and K. Jordaan. Bounds for zeros of Meixner and Kravchuk polynomials. LMS Journal of Computation and Mathematics, 17(1):47–57, 2014. CO-DEN ???? ISSN 1461-1570.

Gonzalez-Jimenez:2014:CRS

[253] Enrique González-Jiménez and Xavier Xarles. On a conjecture of Rudin on squares in arithmetic progressions. LMS Journal of Computation and Mathematics, 17(1):58–76, 2014. CODEN ???? ISSN 1461-1570.

Stoimenow:2014:MGF

[254] A. Stoimenow. Minimal genus and fibering of canonical surfaces via disk decomposition. LMS Journal of Computation and Mathematics, 17(1):77–108, 2014. CODEN ???? ISSN 1461-1570.

Goodwin:2014:CCC

[255] Simon M. Goodwin, Peter Mosch, and Gerhard Röhrle. Calculating conjugacy classes in Sylow p-subgroups of finite Chevalley groups of rank six and seven. LMS Journal of Computation and Mathematics, 17(1):109–122, 2014. CODEN ???? ISSN 1461-1570.

Arora:2014:DPF

[256] Manuel Arora, Gábor Ivanyos, Marek Karpinski, and Nitin Saxena. Deterministic polynomial factoring and association schemes. LMS Journal of Computation and Mathematics, 17(1):123–140, 2014. CODEN ???? ISSN 1461-1570.

Fieker:2014:CGG

[257] Claus Fieker and Jürgen Klüners. Computation of Galois groups of rational polynomials. LMS Journal of Computation and Mathematics, 17(1):141–158, 2014. CODEN ???? ISSN 1461-1570.

Loczi:2014:RFM

[258] Lajos Lóczi and David I. Ketcheson. Rational functions with maximal radius of absolute monotonicity. LMS Journal of Computation and Mathematics, 17(1): 159–205, 2014. CODEN ???? ISSN 1461-1570.

Hamilton:2014:CTO

[259] Thomas Hamilton and David Loeffler. Congruence testing for odd subgroups of the modular group. LMS Journal of Computation and Mathematics, 17(1): 206–208, 2014. CODEN ???? ISSN 1461-1570.

Krasikov:2014:ABA

[260] Ilia Krasikov. Approximations for the Bessel and Airy functions with an explicit error term. LMS Journal of Computation and Mathematics, 17(1):209– 225, 2014. CODEN ???? ISSN 1461-1570.

Bez:2014:NMB

[261] H. E. Bez and N. Bez. A note on magnitude bounds for the mask coefficients of the interpolatory Dubuc–Deslauriers subdivision scheme. LMS Journal of Computation and Mathematics, 17(1): 226–232, 2014. CODEN ???? ISSN 1461-1570.

Moale:2014:NUA

[262] Ionela Moale and Veronika Pillwein. A note on uniform approximation of functions having a double pole. LMS Journal of Computation and Mathematics, 17(1): 233–244, 2014. CODEN ???? ISSN 1461-1570.

Farmer:2014:EFF

[263] David W. Farmer and Nathan C. Ryan.

Evaluating L-functions with few known coefficients. LMS Journal of Computation and Mathematics, 17(1):245–258, 2014. CODEN ???? ISSN 1461-1570.

Siddiqi:2014:SNS

[264] S. S. Siddiqi and M. Younis. A symmetric C^{3} non-stationary subdivision scheme. LMS Journal of Computation and Mathematics, 17(1):259–272, 2014. CODEN ???? ISSN 1461-1570.

Krasikov:2014:BFT

[265] Ilia Krasikov. On the Bessel function in the transition region. LMS Journal of Computation and Mathematics, 17(1): 273–281, 2014. CODEN ???? ISSN 1461-1570.

Dujella:2014:HRE

[266] Andrej Dujella and Juan Carlos Peral.

High-rank elliptic curves with torsion induced by Diophantine triples. LMS Journal of Computation and Mathematics, 17(1):282–288, 2014. CODEN ???? ISSN 1461-1570.

Hichri:2014:BES

[267] Hachem Hichri. On the beta expansion of Salem numbers of degree 8. LMS Journal of Computation and Mathematics, 17(1):289–301, 2014. CODEN ???? ISSN 1461-1570.

Caruso:2014:LAR

[268] Xavier Caruso and David Lubicz. Linear algebra over Z_{p}[[u]] and related rings. LMS Journal of Computation and Mathematics, 17(1):302–344, 2014. CODEN ???? ISSN 1461-1570.

Eick:2014:CMP

[269] B. Eick, M. Kirschmer, and C. Leedham-Green. The constructive membership problem for discrete free subgroups of rank 2 of SL_{2}(R). LMS Journal of Computation and Mathematics, 17(1):345– 359, 2014. CODEN ???? ISSN 1461-1570.

Moehlmann:2014:CMW

[270] G. Moehlmann. Computation of Mordell–Weil bases for ordinary elliptic curves in characteristic two. LMS Journal of Computation and Mathematics, 17(A):1–13, 2014. CODEN ???? ISSN 1461-1570.

Elsenhans:2014:ESR

[271] Andreas-Stephan Elsenhans and Jörg Jahnel. Examples of K3 surfaces with real multiplication. LMS Journal of Computation and Mathematics, 17(A): 14–35, 2014. CODEN ???? ISSN 1461-1570.

Zeng:2014:CGR

[272] Jinxiang Zeng. Computing Galois representations of modular abelian surfaces. LMS Journal of Computation and Mathematics, 17(A):36–48, 2014. CODEN ???? ISSN 1461-1570.

Becker:2014:SAB

[273] Anja Becker, Nicolas Gama, and Antoine Joux. A sieve algorithm based on overlattices. LMS Journal of Computation and Mathematics, 17(A):49–70, 2014. CODEN ???? ISSN 1461-1570.

Chevyrev:2014:CSE

[274] Ilya Chevyrev and Steven D. Galbraith. Constructing supersingular elliptic curves with a given endomorphism ring. LMS Journal of Computation and Mathematics, 17(A):71–91, 2014. CO-DEN ???? ISSN 1461-1570.

Li:2014:ADS

[275] Jianwei Li and Phong Q. Nguyen. Approximating the densest sublattice from Rankin’s inequality. LMS Journal of Computation and Mathematics, 17(A): 92–111, 2014. CODEN ???? ISSN 1461-1570.

Fisher:2014:MMC

[276] Tom Fisher. Minimal models for 6coverings of elliptic curves. LMS Journal of Computation and Mathematics, 17 (A):112–127, 2014. CODEN ???? ISSN 1461-1570.

Lercier:2014:PMS

[277] Reynald Lercier, Christophe Ritzenthaler, Florent Rovetta, and Jeroen Sijsling. Parametrizing the moduli space of curves and applications to smooth plane quartics over finite fields. LMS Journal of Computation and Mathematics, 17(A):128–147, 2014. CODEN ???? ISSN 1461-1570.

Blackburn:2014:DLP

[278] Simon R. Blackburn and Sam Scott. The discrete logarithm problem for exponents of bounded height. LMS Journal of Computation and Mathematics, 17 (A):148–156, 2014. CODEN ???? ISSN 1461-1570.

Costello:2014:CAS

[279] Craig Costello, Alyson Deines-Schartz, Kristin Lauter, and Tonghai Yang. Constructing abelian surfaces for cryptography via Rosenhain invariants. LMS Journal of Computation and Mathematics, 17(A):157–180, 2014. CODEN ???? ISSN 1461-1570.

Bernstein:2014:HEC

[280] Daniel J. Bernstein and Tanja Lange. Hyper-and-elliptic-curve cryptography. LMS Journal of Computation and Mathematics, 17(A):181–202, 2014. CODEN ???? ISSN 1461-1570.

Petit:2014:FRS

[281] Christophe Petit. Finding roots in F_{p^{n}} with the successive resultants algorithm. LMS Journal of Computation and Mathematics, 17(A):203–217, 2014. CODEN ???? ISSN 1461-1570.

Cheng:2014:TBA

[282] Qi Cheng, Daqing Wan, and Jincheng Zhuang. Traps to the BGJT-algorithm for discrete logarithms. LMS Journal of Computation and Mathematics, 17(A): 218–229, 2014. CODEN ???? ISSN 1461-1570.

Barbulescu:2014:MNF

[283] Razvan Barbulescu and Cécile Pierrot. The multiple number field sieve for medium- and high-characteristic finite fields. LMS Journal of Computation and Mathematics, 17(A):230–246, 2014. CO-DEN ???? ISSN 1461-1570.

Takemori:2014:CDV

[284] Sho Takemori. On the computation of the determinant of vector-valued Siegel modular forms. LMS Journal of Computation and Mathematics, 17(A):247–256, 2014. CODEN ???? ISSN 1461-1570.

Harvey:2014:CHW

[285] David Harvey and Andrew V. Sutherland. Computing Hasse–Witt matrices of hyperelliptic curves in average polynomial time. LMS Journal of Computation and Mathematics, 17(A):257–273, 2014. CODEN ???? ISSN 1461-1570.

Caruso:2014:TAP

[286] Xavier Caruso, David Roe, and Tristan Vaccon. Tracking p-adic precision. LMS Journal of Computation and Mathematics, 17(A):274–294, 2014. CODEN ???? ISSN 1461-1570.

Fukuda:2014:CNC

[287] T. Fukuda and K. Komatsu. Class number calculation using Siegel functions. LMS Journal of Computation and Mathematics, 17(A):295–302, 2014. CODEN ???? ISSN 1461-1570.

Kucuksakalli:2014:UGW

[288] Ömer Küçüksakalli. On the units generated by Weierstrass forms. LMS Journal of Computation and Mathematics, 17 (A):303–313, 2014. CODEN ???? ISSN 1461-1570.

Lukas:2014:CQP

[289] David Lukas, Michelle Manes, and Diane Yap. A census of quadratic postcritically finite rational functions defined over Q. LMS Journal of Computation and Mathematics, 17(A):314–329, 2014. CODEN ???? ISSN 1461-1570.

Ryan:2014:NTF

[290] Nathan C. Ryan, Gonzalo Tornaría, and John Voight. Nonvanishing of twists of L-functions attached to Hilbert modular forms. LMS Journal of Computation and Mathematics, 17(A):330–348, 2014. CODEN ???? ISSN 1461-1570.

Fieker:2014:CQR

[291] Claus Fieker and Tommy Hofmann. Computing in quotients of rings of integers. LMS Journal of Computation and Mathematics, 17(A):349–365, 2014. CO-DEN ???? ISSN 1461-1570.

Page:2014:API

[292] A. Page. An algorithm for the principal ideal problem in indefinite quaternion algebras. LMS Journal of Computation and Mathematics, 17(A):366–384, 2014. CODEN ???? ISSN 1461-1570.

Biasse:2014:SCG

[293] Jean-François Biasse and Claus Fieker. Subexponential class group and unit group computation in large degree number fields. LMS Journal of Computation and Mathematics, 17(A):385–403, 2014. CODEN ???? ISSN 1461-1570.

Miller:2014:CNR

[294] John C. Miller. Class numbers of real cyclotomic fields of composite conductor. LMS Journal of Computation and Mathematics, 17(A):404–417, 2014. CODEN ???? ISSN 1461-1570.

Kohel:2014:QIP

[295] David Kohel, Kristin Lauter, Christophe Petit, and Jean-Pierre Tignol. On the quaternion ℓ-isogeny path problem. LMS Journal of Computation and Mathematics, 17(A):418–432, 2014. CODEN ???? ISSN 1461-1570.

Lee:2015:FBP

[296] Edmond W. H. Lee and Wen Ting Zhang. Finite basis problem for semigroups of order six. LMS Journal of Computation and Mathematics, 18:1– 129, 2015. CODEN ???? ISSN 1461-1570.

Almeida:2015:MNF

[297] J. Almeida, J. C. Costa, and M. Zeitoun.

McCammond’s normal forms for free aperiodic semigroups revisited. LMS Journal of Computation and Mathematics, 18:130–147, 2015. CODEN ???? ISSN 1461-1570.

Boyd:2015:SVT

[298] David W. Boyd, Greg Martin, and Mark Thom. Squarefree values of trinomial discriminants. LMS Journal of Computation and Mathematics, 18:148–169, 2015. CODEN ???? ISSN 1461-1570.

Broker:2015:GCJ

[299] Reinier Bröker, Everett W. Howe, Kristin E. Lauter, and Peter Stevenhagen. Genus-2 curves and Jacobians with a given number of points. LMS Journal of Computation and Mathematics, 18:170–197, 2015. CODEN ???? ISSN 1461-1570.

Lubicz:2015:CSI

[300] David Lubicz and Damien Robert. Computing separable isogenies in quasioptimal time. LMS Journal of Computation and Mathematics, 18:198–216, 2015. CODEN ???? ISSN 1461-1570.

Pobegailo:2015:CSC

[301] A. P. Pobegailo. Construction of spline curves on smooth manifolds by action of Lie groups. LMS Journal of Computation and Mathematics, 18(1):217–230, 2015. CODEN ???? ISSN 1461-1570.

Vonk:2015:COF

[302] Jan Vonk. Computing overconvergent forms for small primes. LMS Journal of Computation and Mathematics, 18(1): 250–257, 2015. CODEN ???? ISSN 1461-1570.

Balakrishnan:2015:CIE

[303] Jennifer S. Balakrishnan. Coleman integration for even-degree models of hyperelliptic curves. LMS Journal of Computation and Mathematics, 18(1):258–265, 2015. CODEN ???? ISSN 1461-1570.

Thiel:2015:CCA

[304] U. Thiel. Champ: a Cherednik algebra Magma package. LMS Journal of Computation and Mathematics, 18(1): 266–307, 2015. CODEN ???? ISSN 1461-1570.

Shparlinski:2015:DAE

[305] Igor E. Shparlinski and Andrew V. Sutherland. On the distribution of Atkin and Elkies primes for reductions of elliptic curves on average. LMS Journal of Computation and Mathematics, 18(1): 308–322, 2015. CODEN ???? ISSN 1461-1570.

Ballantyne:2015:MS

[306] John Ballantyne, Chris Bates, and Peter Rowley. The maximal subgroups of E_{7}(2). LMS Journal of Computation and Mathematics, 18(1):323–371, 2015. CODEN ???? ISSN 1461-1570.

Grant:2015:BFD

[307] Timothy J. Grant. Bespoke finite difference schemes that preserve multiple conservation laws. LMS Journal of Computation and Mathematics, 18(1):372–403, 2015. CODEN ???? ISSN 1461-1570.

Corr:2015:NIS

[308] Brian P. Corr, Tomasz Popiel, and Cheryl E. Praeger. Nilpotentindependent sets and estimation in matrix algebras. LMS Journal of Computation and Mathematics, 18(1):404–418, 2015. CODEN ???? ISSN 1461-1570.

Boxall:2015:HPF

[309] John Boxall and David Gruenewald.

Heuristics on pairing-friendly abelian varieties. LMS Journal of Computation and Mathematics, 18(1):419–443, 2015. CODEN ???? ISSN 1461-1570.

Cashen:2015:VGR

[310] Christopher H. Cashen and Jason F.

Manning. Virtual geometricity is rare. LMS Journal of Computation and Mathematics, 18(1):444–455, 2015. CODEN ???? ISSN 1461-1570.

Himstedt:2015:RUC

[311] Frank Himstedt and Felix Noeske. Restricting unipotent characters in special orthogonal groups. LMS Journal of Computation and Mathematics, 18(1): 456–488, 2015. CODEN ???? ISSN 1461-1570.

Duan:2015:SPC

[312] Haibao Duan and Xuezhi Zhao. Schubert presentation of the cohomology ring of flag manifolds. LMS Journal of Computation and Mathematics, 18(1):489– 506, 2015. CODEN ???? ISSN 1461-1570.

Bouyer:2015:ECC

[313] Florian Bouyer and Marco Streng. Examples of CM curves of genus two defined over the reflex field. LMS Journal of Computation and Mathematics, 18(1):

507–538, 2015. CODEN ???? ISSN 1461-1570.

Riza:2015:RKM

[314] Mustafa Riza and Hatice Aktöre. The Runge–Kutta method in geometric multiplicative calculus. LMS Journal of Computation and Mathematics, 18(1): 539–554, 2015. CODEN ???? ISSN 1461-1570.

Couveignes:2015:CFJ

[315] Jean-Marc Couveignes and Tony Ezome. Computing functions on Jacobians and their quotients. LMS Journal of Computation and Mathematics, 18(1):555–577, 2015. CODEN ???? ISSN 1461-1570.

Bruin:2015:HMC

[316] Peter Bruin and Filip Najman. Hyperelliptic modular curves and isogenies of elliptic curves over quadratic fields. LMS Journal of Computation and Mathematics, 18(1):578–602, 2015. CODEN ???? ISSN 1461-1570.

Milio:2015:QLT

[317] Enea Milio. A quasi-linear time algorithm for computing modular polynomials in dimension 2. LMS Journal of Computation and Mathematics, 18(1): 603–632, 2015. CODEN ???? ISSN 1461-1570.

Bennett:2015:MEC

[318] Michael A. Bennett and Amir Ghadermarzi. Mordell’s equation: a classical approach. LMS Journal of Computation and Mathematics, 18(1):633–646, 2015. CODEN ???? ISSN 1461-1570.

Hausen:2015:SPM

[319] Jürgen Hausen and Simon Keicher. A software package for Mori dream spaces.

LMS Journal of Computation and Mathematics, 18(1):647–659, 2015. CODEN ???? ISSN 1461-1570.

Ocalan:2015:ODE

[320] Özkan Öcalan. Oscillation of differential equations with non-monotone retarded arguments. LMS Journal of Computation and Mathematics, 18(1):660–666, 2015. CODEN ???? ISSN 1461-1570.

Wilson:2015:EMC

[321] Robert A. Wilson. Every PSL_{2}(13)

in the Monster contains 13A-elements. LMS Journal of Computation and Mathematics, 18(1):667–674, 2015. CODEN ???? ISSN 1461-1570.

Sorenson:2015:TCI

[322] Jonathan P. Sorenson. Two compact incremental prime sieves. LMS Journal of Computation and Mathematics, 18(1): 675–683, 2015. CODEN ???? ISSN 1461-1570.

Delbourgo:2015:IAC

[323] Daniel Delbourgo and Qin Chao. On λinvariants attached to cyclic cubic number fields. LMS Journal of Computation and Mathematics, 18(1):684–698, 2015. CODEN ???? ISSN 1461-1570.

Bassa:2015:GFD

[324] Alp Bassa, Peter Beelen, and Nhut Nguyen. Good families of Drinfeld modular curves. LMS Journal of Computation and Mathematics, 18(1):699–712, 2015. CODEN ???? ISSN 1461-1570.

Nijmeijer:2015:PRF

[325] M. J. P. Nijmeijer. A parallel rootfinding algorithm. LMS Journal of Computation and Mathematics, 18(1):713– 729, 2015. CODEN ???? ISSN 1461-

Temimi:2015:CIM

[326] H. Temimi and A. R. Ansari. A computational iterative method for solving nonlinear ordinary differential equations. LMS Journal of Computation and Mathematics, 18(1):730–753, 2015. CO-DEN ???? ISSN 1461-1570.

Guglielmetti:2015:CCI

[327] R. Guglielmetti. CoxIter — computing invariants of hyperbolic Coxeter groups. LMS Journal of Computation and Mathematics, 18(1):754–773, 2015. CODEN ???? ISSN 1461-1570.

Cheon:2016:NAD

[328] Jung Hee Cheon and Taechan Kim. A new approach to the discrete logarithm problem with auxiliary inputs. LMS Journal of Computation and Mathematics, 19(1):1–15, January 2016. CODEN ???? ISSN 1461-1570.

Dudek:2016:SSP

[329] Adrian W. Dudek and David J. Platt. On the sum of the square of a prime and a square-free number. LMS Journal of Computation and Mathematics, 19 (1):16–24, January 2016. CODEN ???? ISSN 1461-1570.

Khan:2016:RAE

[330] Naushad Mamode Khan. A robust algorithm for estimating regression and dispersion parameters in non-stationary longitudinally correlated com-Poisson data. LMS Journal of Computation and Mathematics, 19(1):25–36, January 2016. CODEN ???? ISSN 1461-1570.

Platt:2016:ZPS

[331] David J. Platt and Timothy S. Trudgian. Zeroes of partial sums of the zetafunction. LMS Journal of Computation and Mathematics, 19(1):37–41, January 2016. CODEN ???? ISSN 1461-1570.

Rito:2016:CQS

[332] Carlos Rito. Cuspidal quintics and surfaces with p_{g} = 0, K^{2} = 3 and 5torsion. LMS Journal of Computation and Mathematics, 19(1):42–53, January 2016. CODEN ???? ISSN 1461-1570.

Ramos:2016:HNI

[333] Zaqueu Ramos and Aron Simis. Homaloidal nets and ideals of fat points I. LMS Journal of Computation and Mathematics, 19(1):54–77, January 2016. CODEN ???? ISSN 1461-1570.

Kapfer:2016:CCP

[334] Simon Kapfer. Computing cup products in integral cohomology of Hilbert schemes of points on K3 surfaces. LMS Journal of Computation and Mathematics, 19(1):78–97, January 2016. CODEN ???? ISSN 1461-1570.

Chatzarakis:2016:ODE

[335] George E. Chatzarakis and Özkan Öcalan. Oscillation of differential equations with non-monotone retarded arguments. LMS Journal of Computation and Mathematics, 19(1):98–104, January 2016. CODEN ???? ISSN 1461-1570.

Jones:2016:MEC

[336] Andrew Jones. Modular elliptic curves over the field of twelfth roots of unity. LMS Journal of Computation and Mathematics, 19(1):155–174, January 2016. CODEN ???? ISSN 1461-1570.

Lee:2016:FOS

[337] Daniel Lee and Hui-Chun Tien.

A fourth-order seven-point cubature on regular hexagons. LMS Journal of Computation and Mathematics, 19:175–185, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
C9514F36C781D4BCF1BCBCE05E984C11.

Delgado:2016:NSG

[338] M. Delgado, P. A. García-Sánchez, and A. M. Robles-Pérez. Numerical semigroups with a given set of pseudo-Frobenius numbers. LMS Journal of Computation and Mathematics, 19:186–205, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
ADE98085AE61902913CD11BDC225A777.

Arquette:2016:LPR

[339] David M. Arquette and Dursun A. Bulutoglu. The linear programming relaxation permutation symmetry group of an orthogonal array defining integer linear program. LMS Journal of Computation and Mathematics, 19:206–216, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
F30184A13274E530DBA426D22AE1547D.

Fieker:2016:AAA

[340] Claus Fieker and Yinan Zhang. An application of the p-adic analytic class number formula. LMS Journal of Computation and Mathematics, 19:217–228, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
5AF6C8AA639096F737C10056CA814DB9.

Bray:2016:FBM

[341] John N. Bray, Richard A. Parker, and Robert A. Wilson. Finding 47: 23 in the Baby Monster. LMS Journal of Computation and Mathematics, 19:229–234, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
364EB017BA9562E79FFBDEBF8312CA5D.

Stewart:2016:MME

[342] David I. Stewart. On the minimal modules for exceptional Lie algebras: Jordan blocks and stabilizers. LMS Journal of Computation and Mathematics, 19:235–258, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
5C0D84132C8C8B9B0312A6D73A802241.

Bucur:2016:LTC

[343] Alina Bucur, Anne-Maria Ernvall-Hytönen, Almasa Odzak, and Lejla Smajlović. On a Li-type criterion for zero-free regions of certain Dirichlet series with real coefficients. LMS Journal of Computation and Mathematics, 19:259–280, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
9D6498CFDB802707E1A0EDFFE5B81C34.

Lochowski:2016:DLT

[344] Rafal Marcin Lochowski. On the double Laplace transform of the truncated variation of a Brownian motion with drift. LMS Journal of Computation and Mathematics, 19:281–292, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
C1F4B21705E4BC3A4A0BF6BA7A2FBF1D.

Wang:2016:ASC

[345] Xiuhua Wang and Jisheng Kou. Analysis of semilocal convergence for ameliorated super-Halley methods with less computation for inversion. LMS Journal of Computation and Mathematics, 19:293–302, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
BAA0A93068203107B0C85B0D874E1EBE.

Himstedt:2016:CSS

[346] Frank Himstedt, Tung Le, and Kay Magaard. On the characters of the Sylow p-subgroups of untwisted Chevalley groups Y_{n}(p^{a}). LMS Journal of Computation and Mathematics, 19(2):303–359, 2016. CODEN ???? ISSN 1461-1570.

Kedlaya:2016:CZF

[347] Kiran S. Kedlaya and Andrew V. Sutherland. A census of zeta functions of quartic K3 surfaces over F_{2}. LMS Journal of Computation and Mathematics, 19(A):1–11, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
514C89A1FE39BEE2565518DA51DB3B5C.

Elsenhans:2016:PCK

[348] Andreas-Stephan Elsenhans and Jörg Jahnel. Point counting on K3 surfaces and an application concerning real and complex multiplication. LMS Journal of Computation and Mathematics, 19(A):12–28, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
BB5A2410DBE981DDC9F9B5E9BEF54760.

Kumar:2016:RMT

[349] Abhinav Kumar and Ronen E. Mukamel. Real multiplication through explicit correspondences. LMS Journal of Computation and Mathematics, 19 (A):29–42, 2016. CODEN ???? ISSN 1461-1570. URL https:/

/www.cambridge.org/core/product/
1478E13EEB8BAE631C6D1CA10809A86B.

vanVredendaal:2016:RMM

[350] Christine van Vredendaal. Reduced memory meet-in-the-middle attack against the NTRU private key. LMS Journal of Computation and Mathematics, 19(A):43–57, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
2FD6898DA25DD88B007F12A56421BA73.

Galbraith:2016:AAC

[351] Steven D. Galbraith, Shishay W. Gebregiyorgis, and Sean Murphy. Algorithms for the approximate common divisor problem. LMS Journal of Computation and Mathematics, 19(A):58–72, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
2E0B1F3C002E4AF310B0AD368B1F80E7.

Quertier:2016:EHP

[352] Tony Quertier. Effective Hasse principle for the intersection of two quadrics. LMS Journal of Computation and Mathematics, 19(A):73–82, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
FCEAC4AE2C22F81B1A82A528F6DEB7BD.

Heer:2016:JEI

[353] Henriette Heer, Gary McGuire, and Oisín Robinson. JKL-ECM: an implementation of ECM using Hessian curves. LMS Journal of Computation and Mathematics, 19(A):83–99, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
30EE6BE8CCF8A6F0592B3B536B08326E.

Fisher:2016:VEO

[354] Tom Fisher. Visualizing elements of order 7 in the Tate–Shafarevich group of an elliptic curve. LMS Journal of Computation and Mathematics, 19(A):100–114, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
DA5E76EABC9AE06B22E5FE01A5AC8482.

Morain:2016:CCC

[355] François Morain, Charlotte Scribot, and Benjamin Smith. Computing cardinalities of Q-curve reductions over finite fields. LMS Journal of Computation and Mathematics, 19(A):115–129, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
3BA9531DF1A0887DC7FB71914EB0095C.

Castryck:2016:EDR

[356] Wouter Castryck, Ilia Iliashenko, and Frederik Vercauteren. On error distributions in ring-based LWE. LMS Journal of Computation and Mathematics, 19(A):130–145, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
C852C8D1C2F84DA398EA07D2B2F3E4CB.

Bai:2016:TLS

[357] Shi Bai, Thijs Laarhoven, and Damien Stehlé. Tuple lattice sieving. LMS Journal of Computation and Mathematics, 19(A):146–162, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
C1CE6384DEC54330AEFB2A4D38190094.

Labrande:2016:CTF

[358] Hugo Labrande and Emmanuel Thomé. Computing theta functions in quasilinear time in genus two and above. LMS Journal of Computation and Mathematics, 19(A):163–177, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
2EE73D9E67B049C98C369F681F28741B.

Bockle:2016:DAM

[359] Gebhard Böckle and Damián Gvirtz.

Division algebras and maximal orders for given invariants. LMS Journal of Computation and Mathematics, 19(A):178–195, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
7A9F4731660F6140FDF90662210534CE.

Kelley:2016:RSP

[360] Zander Kelley. Roots of sparse polynomials over a finite field. LMS Journal of Computation and Mathematics, 19(A):196–204, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
0B4BB0B58AF86C34B8D12395335A3958.

Ryan:2016:CJF

[361] Nathan C. Ryan, Nicolás Sirolli, Nils-Peter Skoruppa, and Gonzalo Tornaría. Computing Jacobi forms. LMS Journal of Computation and Mathematics, 19(A):205–219, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
B9816899D93044A71CF095492887893D.

Harvey:2016:CSG

[362] David Harvey, Maike Massierer, and Andrew V. Sutherland. Computing L-series of geometrically hyperelliptic curves of genus three. LMS Journal of Computation and Mathematics, 19(A):220–234, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
0C92F653663C97385F1B2780A120A804.

Booker:2016:DGC

[363] Andrew R. Booker, Jeroen Sijsling, Andrew V. Sutherland, John Voight, and Dan Yasaki. A database of genus-2 curves over the rational numbers. LMS Journal of Computation and Mathematics, 19(A):235–254, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
5D8F4D220534D6EEEA18133590545035.

Cheon:2016:ANP

[364] Jung Hee Cheon, Jinhyuck Jeong, and Changmin Lee. An algorithm for NTRU problems and cryptanalysis of the GGH multilinear map without a low-level encoding of zero. LMS Journal of Computation and Mathematics, 19(A):255–266, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
230ECFEEE6AF4D8027FF3E13998D560C.

DeFeo:2016:EIQ

[365] Luca De Feo, Cyril Hugounenq, Jérôme Plût, and Éric Schost. Explicit isogenies in quadratic time in any characteristic. LMS Journal of Computation and Mathematics, 19(A):267–282, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
E22F1986410DD4A0CC4B1645C0E7538F.

Balakrishnan:2016:CGH

[366] Jennifer S. Balakrishnan, Sorina Ionica, Kristin Lauter, and Christelle Vincent. Constructing genus-3 hyperelliptic Jacobians with CM. LMS Journal of Computation and Mathematics, 19(A):283–300, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
1F6A12321A3E445E69BC648DF24DBC10.

Shieh:2016:CTA

[367] Yih-Dar Shieh. Character theory approach to Sato–Tate groups. LMS Journal of Computation and Mathematics, 19(A):301–314, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
34497E7C4A3FB27857A14FB82621B8CE.

Gelin:2016:RNF

[368] Alexandre Gélin and Antoine Joux.

Reducing number field defining polynomials: an application to class group computations. LMS Journal of Computation and Mathematics, 19(A):315–331, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
8A2022D2F9DC244A606F4EB264EDF2D9.

Gaudry:2016:CRN

[369] Pierrick Gaudry, Laurent Grémy, and Marion Videau. Collecting relations for the number field sieve in GF(p^{6}). LMS Journal of Computation and Mathematics, 19(A):332–350, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
7CA4DA868670B359E88C9663E4B52F92.

Balakrishnan:2016:DEC

[370] Jennifer S. Balakrishnan, Wei Ho, Nathan Kaplan, Simon Spicer, William Stein, and James Weigandt. Databases of elliptic curves ordered by height and distributions of Selmer groups and ranks. LMS Journal of Computation and Mathematics, 19(A):351–370, 2016. CODEN ???? ISSN 1461-1570. URL

https://www.cambridge.org/core/
product/E245E7A5300FEA32E48E89A71C746930.

Biasse:2016:FHA

[371] Jean-François Biasse, Claus Fieker, and Michael J. Jacobson. Fast heuristic algorithms for computing relations in the class group of a quadratic order, with applications to isogeny evaluation. LMS Journal of Computation and Mathematics, 19(A):371–390, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
7FEE0CD0D530B9B3F2EAD1B71340BBA9.

Muller:2016:CCH

[372] J. Steffen Müller and Michael Stoll. Computing canonical heights on elliptic curves in quasi-linear time. LMS Journal of Computation and Mathematics, 19(A):391–405, 2016. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
B0FC77ECD1B5A0F4FD8A848D3671EDF4.

Lv:2017:TTO

[373] Xueqin Lv and Jianfang Gao. Treatment for third-order nonlinear differential equations based on the Adomian decomposition method. LMS Journal of Computation and Mathematics, 20(1):1–10, 2017. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
D35BF38BABE089B43D5AEB1652C4AE16.

Shah:2017:GSB

[374] Kamal Shah, Hammad Khalil, and Rahmat Ali Khan. A generalized scheme based on shifted Jacobi polynomials for numerical simulation of coupled systems of multi-term fractional-order partial differential equations. LMS Journal of Computation and Mathematics, 20(1):11–29, 2017. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
36ACCE9EA815D993F53CF6000D8DA78E.

Basraoui:2017:SEC

[375] Abdelkrim El Basraoui and John McKay. The Schwarzian equation for completely replicable functions. LMS Journal of Computation and Mathematics, 20(1):30–52, 2017. CODEN ???? ISSN 1461-1570. URL https:

//www.cambridge.org/core/product/
6ECB7A9C89D3B939CF1BA1D5EBDB8BDB.